### v2

# ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection

Abstract
Ensuring neural networks adhere to domain-specific constraints is crucial for addressing safety and ethical concerns while also enhancing prediction accuracy. Despite the nonlinear nature of most real-world tasks, existing methods are predominantly limited to affine or convex constraints. We introduce ENFORCE, a neural network architecture that guarantees predictions to satisfy nonlinear constraints exactly. ENFORCE is trained with standard unconstrained gradient-based optimizers (e.g., Adam) and leverages autodifferentiation and local neural projections to enforce any $\mathcal{C}^{1}$ constraint to arbitrary tolerance $\epsilon$. We build an adaptive-depth neural projection (AdaNP) module that dynamically adjusts its complexity to suit the specific problem and the required tolerance levels. ENFORCE guarantees satisfaction of equality constraints that are nonlinear in both inputs and outputs of the neural network with minimal (and adjustable) computational cost.
Keywords Constrained learning  $\cdot$ Hard-constrained neural networks  $\cdot$ Trustworthy AI  $\cdot$ Physics-informed machine learning
1 Introduction
Neural networks (NNs) are the backbone of many recent advancements in artificial intelligence (AI), excelling in tasks such as natural language processing, image analysis, and scientific discovery due to their modularity, simplicity, and strong generalization capabilities. However, their ability falls short when strict adherence to domain-specific constraints is required. Depending on the task, prior knowledge and rules about the system (e.g., from physics, safety, or ethics) are often available and are typically leveraged by humans in decision-making processes. In contrast, traditional data-driven methods such as NNs rely solely on data. Thus, trained NNs may be accurate on a training and test data set but still may not satisfy constraints exactly, leading to inconsistent predictions. This limitation not only generates substantial skepticism, hindering their adoption in real-world applications but can potentially result in erroneous or unfeasible outcomes when NNs are applied to decision-making tasks. Moreover, when prior domain knowledge is available and can be expressed as analytical expressions such as equations, ensuring that NNs adhere to this information is crucial to avoid the suboptimal utilization of valuable insights.
Enforcing strict constraints in NNs is a promising area of research for many fields. For example, in AI for Science, integrating first-principle laws ensures physically consistent models enabling insightful scientific discovery [1, 2]. More broadly, constraining neural network predictions can have a transformative impact in domains where strict adherence to critical requirements is essential. This includes applications such as safety-critical systems (e.g., self-driving vehicles [3], healthcare [4]), fairness and bias mitigation [5, 6], and compliance with regulatory standards (e.g., in finance [7]). Additionally, with the rise of generative AI (GenAI), enforcing constraints on generation processes could mitigate potential risks and ensure trustworthy adherence to given criteria. Constraining the neural network output to adhere to strict rules is also beneficial to tackle traditional machine learning challenges, such as overfitting and underfitting in data-scarce regimes [8]. It is also plausible that such constraints could enable some level of extrapolation in the presence of incomplete data or in holes of convex hulls [9, 10]. Hence, the combination of enhanced accuracy and adherence to pre-existing domain knowledge could enable the safe use of NNs as surrogate models in computationally expensive applications [11].
Enforcing constraints in NNs is not an easy task. Early approaches often relied on incorporating penalty terms into the loss function to reduce violations of physical laws [12]. Yet, these penalty-based methods offer no guarantees of constraint satisfaction (soft-constrained). In contrast, other methods aim to ensure strict adherence to analytical constraints by design (hard-constrained). For instance, one can use activation functions such as sigmoids or ReLU to bound outputs. To enforce analytical constraints beyond bounds, recent studies incorporate correction layers into NNs to project or complete predictions, ensuring they lie within the feasible region. Recent research has developed efficient methods to enforce analytical constraints defined by affine relationships between input and output variables or by convex regions (See Section 2). However, many real applications such as in science or sociology are inherently governed by nonlinear relationships between the involved variables [13, 14]. Existing approaches for handling nonlinear constraints predominantly rely on external root-finding algorithms or constrained optimization solvers (Section 2.2). These methods introduce significant computational overhead and complicate model construction, thereby compromising the modularity and flexibility typically associated with NNs.
In this paper, we propose ENFORCE, a neural network architecture that ensures predictions strictly satisfy nonlinear constraints representing domain knowledge or critical requirements. ENFORCE is trained using standard unconstrained optimization techniques and leverages an adaptive-depth projection module to enforce constraints by construction without relying on external solvers.
2 Related work
This section reviews existing approaches to enforce constraints in NNs. According to the focus of this paper, a more in-depth analysis will be conducted on existing methods for enforcing algebraic equality constraints exactly.
2.1 Soft-constrained neural networks
One of the earliest approaches to embedding domain knowledge into NNs involves the use of soft constraints. Soft constraints are incorporated as penalty terms appended to the loss function, penalizing residuals of algebraic [15, 16] or differential equations underlying the system [17]. Physics-informed neural networks (PINNs) [12] represent a widely used framework designed to solve partial differential equations (PDEs) with NNs by employing soft constraints and collocation points. Although the soft-constrained approach places no restrictions on the complexity of the constraints, it has the drawback of not guaranteeing strict adherence to them. Furthermore, increasing the complexity of the loss function – especially when the different terms vary in nature or scale – can degrade the optimization performance of the neural network, often resulting in suboptimal accuracy [18, 19].
2.2 Hard-constrained neural networks
Hard-constrained NNs refer to methodological approaches ensuring that neural network predictions strictly adhere to analytical constraints. These constraints, explicitly encoded within the architecture, act as inductive biases, guiding the learning process toward compliance with domain knowledge or restrictions [20]. Architectures such as convolutional neural networks (CNNs) [21] and graph neural networks (GNNs) [22, 23] encode inductive biases by guaranteeing invariance with respect to patterns and symmetries. Simple analytical constraints can be enforced using differentiable functions, such as sigmoids or ReLU for output bounding and softmax for simplex constraints (non-negativity and normalization). Recent literature includes significant contributions for enforcing analytical inequality constraints, such as convex polytopes and convex sets more generally [24, 25, 26, 27, 28]. One can also constrain the neural network to guarantee specific functional characteristics, such as being 1-Lipschitz [29] or Lyapunov stable [30], nevertheless, this is not a core objective of this study. For a broad and recent review on hard-constrained NNs, the reader is also referred to [8].
Since this paper focuses on analytical equality constraints, the following literature review considers existing methods for this specific case.
Projection methods
Many methods for encoding hard equality constraints utilize projection techniques, which correct preliminary neural network predictions by appending a non-trainable layer to the output. Projections can be formulated as optimization problems (e.g., distance minimization) or derived from geometric principles. For example, in [31] NN predictions of physical systems governed by PDEs are projected to ensure solutions satisfy the finite difference discretization of the underlying linear PDEs. A more general approach is the KKT-hPINN [32], which enforces linear equality constraints of the form $Ax+By=b$, though its applicability is limited to affine equations. Recently, HardNet [8] was introduced to enforce equality and inequality constraints affine in the output, without input restrictions, via a closed-form projection step. HardNet-Cvx [8] extends this approach to more general convex functions in the inputs, leveraging differentiable optimization layers supported by optimization solvers [33].
Predict-and-complete
Alternatively, the neural network can predict a partial set of output variables, $y_{P}\in\mathbb{R}^{N_{O}-N_{C}}$, and complete the prediction by solving the system of constraints based on this partial output. This approach ensures that the constraints are always satisfied. For instance, Beucler et al. introduced this concept to simulate physical systems such as climate modeling [34]. However, when the constraints are not explicitly defined, solving the system requires a root-finding solver. Similar approaches have been proposed within the hybrid modeling community, particularly in the serial configuration, where a fully data-driven method is used to predict certain inputs to a mechanistic model [35]. While studies like DC3 [25] have developed efficient backpropagation techniques, scenarios involving implicit nonlinear constraints can be computationally expensive to tackle with predict-and-complete methods. Moreover, the predict-and-complete approach can reduce regression accuracy by hindering the representation capabilities of NNs [34].
Constrained optimization
One of the most popular fields of application of constrained learning is for solving (parametric) constrained optimization problems using NNs [36]. OptNet [37] is an optimization layer developed to solve quadratic programs. Angrawal et al. [33] expands the methodology to convex programs. They develop efficient differentiation techniques through such layers. However, the forward pass always requires the solution of a constrained optimization problem. Recently, Mukherjee and Bhattacharyya [38] brute-forced the constrained learning paradigm by training a neural network using an NLP solver such as IPOPT [39] instead of standard unconstrained optimization algorithms. Also, NNs have been formulated as constrained optimization problems and solved using deterministic global solvers [11]. However, these approaches pose severe limitations in terms of NNs and dataset size.
Other methods
Other methods have been proposed for constrained learning in NNs, mostly considering affine or convex regions. Many of them consider constraints only dependent on the input of the neural network [10, 27, 40, 41], others design strategies to include the dependence on both inputs and outputs [28, 42]. Recently, contributions to enforce general logic and linear constraints have been proposed by the neuro-symbolic AI community, developing constraining layers using logic programming to improve multi-label classification and synthetic data generation [43, 44].
However, to the best of our knowledge, there is no method designed to enforce any nonlinear equality constraints involving both the input and the output of the neural network by encoding them in the architecture and without relying on external solvers.
3 Preliminaries
Problem statement
Given a dataset $(x_{i},y_{i})_{i=1,...,N}$, we aim to train a neural network $f_{\theta}$ with parameters $\theta$ to approximate the underlying relationships while satisfying a set of known algebraic equality constraints $c(x,y)=0$. In general, $c$ can be any nonlinear function in the input $x$ and output $y$ of the neural network, incorporating domain knowledge or specifying critical requirements.
Assumption 1.
i) The constraints $c(x,y)=0$ are feasible, ii) $N_{C}<N_{O}$, where $N_{C}$ is the number of equality constraints and $N_{O}$ is the output dimensionality of the neural network, i.e., there are available degrees of freedom to learn.
One way to enforce the neural network prediction $\hat{y}$ to satisfy the constraints is to project it onto the feasible hypersurface defined by $c(x,y)=0$. The projection can be defined as a mathematical program:
$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{T}W\|y-\hat{y}\|$ (1)
s.t. $\quad c(x,y)=0$
[…]
Quadratic projection
When $c(x,y)$ is an affine function in the neural network input and output (i.e., $c(x,y)=Ax+By-b=0$), then the problem results in a quadratic program (QP) and a closed-form analytical solution is available for the KKT conditions. Based on such observation, Chen et al. proposed KKT-hPINN by stacking a fixed non-trainable projection layer on top of a neural network, to guarantee the prediction adheres to given affine functions [32]. However, the method is restricted to linear constraint as a closed-form is not available when $c$ is nonlinear.
We note that a closed-form is still available when generalizing to any function in the input $x$, as the projection operation is still a QP. We observe that, assuming an affine dependence on $y$ such that $c=C(x)y-v(x)-b=0$, $C(x)$ and $v(x)$ play the roles of linear coefficients and translation, respectively, in a local sense. That is, given a input $x_{i}$ and a prediction $\hat{y}_{i}$, then any function of $x_{i}$ is a constant term with respect to the optimization problem in Eq. 1, which becomes again a QP.
Enforcing affine functions is not new and is achieved also through other techniques [8, 40]. However, relaxing the assumption to allow for any nonlinear constraints of both the input and output of the neural network commonly results in decreased computational efficiency, as typically requires the use of constrained optimization techniques [38] or nonlinear solvers such as Newton’s methods [25].
4 Nonlinear constrained learning
We present ENFORCE, a framework designed for general and efficient nonlinear constrained learning. The method employs a computationally cheap adaptive neural projection and has no restriction on the nonlinearity of $\mathcal{C}^{1}$ constraints involving both the input and output vectors of the neural network. At inference time, the model predictions satisfy the constraints to arbitrary tolerance $\epsilon$.
4.1 AdaNP: Adaptive-depth neural projection
We reformulate the NLP in Eq. 1 and exploit the efficiency of quadratic projections to generalize the methodology to any nonlinear constraint. Assuming $c$ is of class $\mathcal{C}^{1}$, we use first-order Taylor expansion to locally linearize the constraints around the neural network input $x_{0}$ and prediction $\hat{y}$:
$c(x,y)\simeq c(x_{0},\hat{y})+\left.J_{x}c\right|_{x_{0},\hat{y}}(x-x_{0})+\left.J_{y}c\right|_{x_{0},\hat{y}}(y-\hat{y}),$ (2)
where $J_{x}c$ and $J_{y}c$ are the Jacobian matrices with respect to the variable $x$ and $y$, respectively. Since the neural network input is fixed for a given sample, the approximation is exact in $x$, hence, $x=x_{0}$. Considering orthogonal projection for notation simplicity, the nonlinear optimization problem in Eq. 1 is locally approximated by a (linearly constrained) QP:
$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{2}$ (3)
s.t. $\quad c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
Definition 1.
Given an input $x\in\mathbb{R}^{N_{I}}$ to a neural network $f_{\theta}$, its prediction $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N_{O}}$, and a set of constraints $c\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N{{}_{C}}})$, with $N_{C}<N_{O}$, we define an operator $\mathcal{P}$ such that $\tilde{y}=\mathcal{P}(\hat{y})$ is the solution to the linearized quadratic program in Eq. 3, in the domain $\Omega$ where the constraints are defined.
In particular, $\tilde{y}=B^{*}\hat{y}+v^{*}$, with $B^{*}=I-B^{T}(BB^{T})^{-1}B$ and $v^{*}=B^{T}(BB^{T})^{-1}v$, where $I\in\mathbb{R}^{N_{O}\times N_{O}}$ is the identity matrix, and:
$B$ $=\left.J_{y}c\right|_{x,\hat{y}}$
$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})$
Since a closed-form expression for the operator $\mathcal{P}$ is derived (Appendix A.1), we can define a differentiable neural projection (NP) layer representing the operator $\mathcal{P}$. The forward and backward passes of an NP layer are computationally cheap (more details on implementation and computational cost are given in Appendix A.2). However, the operator $\mathcal{P}$ projects the neural network prediction onto a linear approximation of the nonlinear constraints (i.e., the tangent hyperplane). The error that we introduce is proportional to the projection displacement $\ell_{p}=||\tilde{y}-\hat{y}||$. From this consideration, it follows that (1) the error is mitigated as the projection displacement is small, i.e., the neural network prediction is sufficiently accurate, and (2) a single NP layer may not ensure exact adherence to nonlinear constraints. Nevertheless, it is worth noting that a single NP layer guarantees strict satisfaction of equality constraints that are affine in $y$ and nonlinear in $x$, i.e., it efficiently enforces constraint classes considered in similar recent works [32, 8].
To address the challenge of ensuring exact adherence to nonlinear constraints, as highlighted in the second consideration above, we propose AdaNP: an adaptive-depth neural projection composition to enforce nonlinear constraint satisfaction to arbitrary tolerance $\epsilon$.
Definition 2.
Given an operator $\mathcal{P}$ as defined in Def. 1, AdaNP is a composition of $n$ operators $\mathcal{P}$, such that:
$\text{AdaNP}=\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n}$
Then, given an arbitrarily small scalar $\epsilon$ and $n\in\mathbb{N}$ such that $|c(x,\tilde{y}_{n})|<\epsilon$, the feasible prediction $\tilde{y}_{n}$ is computed as:
$\tilde{y}_{n}=(\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n})(\hat{y})$
AdaNP is a differentiable stack of $n$-NP layers that can be composed on every neural network backbone. The depth $n$ adjusts adaptively depending on the nonlinearities and the specified tolerance. Accurate NNs typically result in shallower AdaNP modules, since the projection displacement $\ell_{p}$ is related to the loss $\ell_{d}=||\hat{y}-y^{*}||$, where $y^{*}$ is the ground truth output. This introduces a trade-off between the complexity of the backbone and the required depth of AdaNP to satisfy the specified tolerance criteria. It is worth mentioning that a similar concept of iterative neural projections, although employed for a different application and objective, is also utilized in [45].
Analogy with Sequential Quadratic Programming
AdaNP can also be seen as an iterative method that recursively improves the solution of a linearized nonlinear program. Here, we notice the similarity to sequential quadratic programming (SQP) techniques. Specifically, AdaNP is a simple case of SQP method for which the objective function is naturally quadratic while the nonlinear constraints are linearized (in contrast to full SQP, in which the objective function is quadratically approximated). This observation could potentially allow to analyze the convergence of the method using SQP theory [46, 47].
Deviation from Newton’s method
While the KKT conditions for an NLP (Eq. 1) can be more generally solved using Newton’s methods, our method circumvents the computational overhead associated with calculating the Hessian matrix of the constraints (Appendix A.1).
4.2 Architecture
The architecture of ENFORCE (Fig. 1) is composed of (1) a backbone neural network, that can be of any kind and complexity, and (2) an AdaNP module. The depth of AdaNP depends on the backbone performance and specified tolerance. Indeed, AdaNP can be purposely tuned to increase training efficiency (Section 4.4). A single NP layer is composed of two steps: (1) automatic differentiation and (2) local neural projection.
Figure 1: ENFORCE consists of a backbone neural network and an adaptive neural projection (AdaNP) module. The backbone network can be of every kind, such as fully connected, convolutional, or transformer architecture. AdaNP includes an adaptive number of neural projection (NP) layers, each composed of an auto-differentiation and a local projection step.
Exact Jacobian computation
To compute the Jacobian of the constraint system, we leverage automatic differentiation available in most deep learning libraries [48, 49, 50]. Computing the local Jacobian $J_{y}c|_{x,\hat{y}}$ is computationally inexpensive, as it requires propagating derivatives only through the constraints and does not involve the neural network backbone. Furthermore, its computation can be efficiently parallelized on GPU.
Local neural projection
The neural projection defined by the operator $\mathcal{P}$ in Def. 1 depends on individual input-prediction instances. Thus, the projection is locally defined in the neighborhood of ($x_{i},\hat{y}_{i}$). We parallelize the computation of local neural projections by building a rank-3 tensor $\mathbf{B}^{*}$ and a rank-2 tensor $\mathbf{v}^{*}$ (Appendix A.2). Thus, we reduce the cost of an NP layer from $\mathcal{O}(BS\times N_{C}^{3})$ to $\mathcal{O}(N_{C}^{3})$, allowing effective training with stochastic gradient descent techniques with virtually no limitation on the batch dimension ($BS$). For modern hardware, handling output dimensionalities of $N_{C}<10^{3}$ leads to minimal computational cost.
4.3 Implications of orthogonal projections
The set of constraints $c(x,y)=0$ describes an infinite-wide feasible region (i.e., a hypersurface) where the constraints are defined. Hence, one could ask whether the projections are unique and thus suitable for guiding the learning process. To this end, theoretical insights can be derived by analyzing the case of orthogonal projections, i.e., $W=I$ (Eq. 1).
(a) Affine constraint
(b) Nonlinear constraint
Figure 2: Orthogonal projection on the feasible region when $y\in\mathbb{R}^{2}$. Orthogonal projections on affine constraints are unique and always lead to enhanced predictions. When the constraint is nonlinear, multiple orthogonal projections are possible, not necessarily improving the prediction.
Affine constraints
Considering constraints that are affine in $y$, the feasible space is a hyperplane defined in $\mathbb{R}^{N_{O}}$. Under those circumstances, it can be proven that there exists a unique orthogonal projection defining $\tilde{y}$ from a point $\hat{y}\in\mathbb{R}^{N_{O}}$ such that $c(\tilde{y})=0$. More interestingly, if $y^{*}$ is the ground truth, then $\forall\hat{y},y^{*},\,||y^{*}-\tilde{y}||<||y^{*}-\hat{y}||$, that is, the projected prediction is always a better prediction than the original one (Fig. 2(a)). In such a condition, constrained learning effectively guides the learning process toward the ground truth. In particular, the loss is expected to decrease compared to the unconstrained case from the very start of the training process.
Nonlinear constraints
The same considerations do not hold for nonlinear hypersurfaces. Firstly, there could be multiple points on the hypersurface where the normal vectors point toward the prediction $\hat{y}$. Secondly, the closest orthogonal projection can potentially represent a worse prediction, since the distance $||y^{*}-\tilde{y}||$ is not ensured to be shorter than $||y^{*}-\hat{y}||$ (see Fig. 2(b)). Then, the projection step could potentially conflict with the training step, preventing the training process itself from converging. Therefore, we can not conclude that orthogonal projections onto a nonlinear feasible region always guide the learning process efficiently.
4.4 Training ENFORCE
We train ENFORCE using standard unconstrained gradient descent methods (i.e., Adam). We develop and use a constrained learning methodology using AdaNP to guide the neural network training to convergence, supported by the theoretical implications of orthogonal projections described above (Section 4.3).
Linearized constraints and training strategy
ENFORCE does not project the predictions directly on the nonlinear feasible region but on a locally linearized counterpart. Thus, we can assume the existence of a unique closest orthogonal projection. However, the quality of the latter, i.e., whether the projection really improves the prediction, is largely influenced by the constraint geometry curvature. Moreover, the initialization of the neural network plays a crucial role in constrained learning using ENFORCE. Overall, it is difficult to prove a priori whether constrained learning can be beneficial during the initial phase of training.
To overcome this limitation, we develop a strategy to ensure effective constrained learning. Instead of building a static ENFORCE model, we activate AdaNP only when its application enhances prediction accuracy (i.e., the loss decreases). This strategy can be interpreted as an unconstrained pre-training; however, it is automatically incorporated in ENFORCE. We have demonstrated that this approach accelerates the training process and leads to better optima compared to fully unconstrained training.
Loss function
We further control the constrained learning process by minimizing the projection displacement, i.e., the distance between the preliminary NN prediction $\hat{y}$ and its projection $\tilde{y}=\text{AdaNP}(\hat{y})$. This aims to (1) ensure minor linearization error ($\epsilon_{L}\sim\Delta y$) and (2) prevent the neural network from learning alternative functions whose projections fall within the neighborhood of the desired functions. The latter is suggested to reduce reliance on AdaNP for accurate predictions, thereby lowering the computational cost during inference (i.e., by decreasing the depth of AdaNP).
We control the projection displacement by applying soft constraints, adding an extra penalty term to the loss function:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\tilde{y}_{i}||^{2}+\frac{\lambda}{N}\sum_{i=1}^{N}||\hat{y}_{i}-\tilde{y}_{i}||^{2}$ (4)
Here, the first term represents the conventional mean squared error between the projected prediction $\tilde{\mathbf{y}}_{i}$ and the ground truth $\mathbf{y}_{i}$, while the second term quantifies the projection displacement, targeted at controlling and minimizing its magnitude, by acting as a penalty. The two terms are balanced by a scalar weight $\lambda$.
5 Experiments and discussion
We demonstrate our proposed method on an illustrative two-dimensional example. We aim to fit the oscillating function $y:\mathbb{R}\rightarrow\mathbb{R}^{2}$:
$y_{1}$ $=2\text{sin}(fx)$ (5)
$y_{2}$ $=-\text{sin}^{2}(fx)-x^{2}\text{,}$
where $x\in\mathbb{R}$ is the unidimensional independent variable (input) and the scalar $f$ is the frequency. Notably, the system is implicitly linked by a nonlinear constraint $c(x,y_{1},y_{2})=(0.5y_{1})^{2}+x^{2}+y_{2}$, involving both input and output variables. We train an ENFORCE model consisting of a 64-neuron 1-hidden layer fully connected ReLU neural network as a backbone and an AdaNP module to force the predictions to satisfy the constraint. We experiment with different constraint tolerances ($\epsilon_{T}$) and loss weighting factors ($\lambda$) during training to heuristically study their influence. At inference, the depth of AdaNP is adapted to guarantee a specified tolerance $\epsilon$, e.g., absolute average or maximum residual across a batch of prediction. In this case study, we set the average absolute residual to be below $\epsilon_{I}=10^{-6}$ at inference time. To verify the interpolation capabilities, we sample 100 training data points from a uniform distribution in $x=[-2,2]$ and 1000 test points in the same domain. Every run is repeated 5 times using different initialization seeds.
Accuracy and constraint guarantee
One of the major concerns in constrained learning is related to keeping the expressivity and approximation capabilities of NNs [51, 52] while enforcing constraints to be satisfied. Here, we evaluate the regression capabilities and robustness to spectral biases [53] of ENFORCE on oscillating functions ($f=5$) while assessing the average and maximum constraint deviation (i.e., residual). We compare the method with a traditional (unconstrained) multilayer perceptron (MLP) and a soft-constrained neural network sharing the same architecture. In the soft-constrained neural network, the constraint residual is added as an additional loss term balanced with a scalar weight ($\lambda_{y}$). We report here the outcomes of the best-performing soft-constrained network, with $\lambda_{y}=1$. We train all the NNs for 50,000 epochs, using Adam [54] optimizer with a learning rate of $0.001$.
The main results are summarized in Table 1. ENFORCE outperforms the soft-constrained NN and the MLP by effectively minimizing the nonlinear constraint residual, guaranteeing arbitrary satisfaction with minor computational costs (Fig. 3(c)). To quantify the relative impact, we normalized the residuals by scaling them with respect to the absolute maximum value of the variables within the considered range ($\max(|x|,|y_{1}|,|y_{2}|)$). It is worth noting how overall good performance (e.g., in NN) does not guarantee compliance with underlying model constraints, with deviations above 17%. By contrast, ENFORCE demonstrates a worst-case residual below 0.0001% (Table 1). The inference time for a batch of 1000 samples is 6 ms longer when using ENFORCE compared to an MLP. This amount should be regarded as additive, not multiplicative (e.g., 4x relative to the MLP). Indeed, the computational complexity is entirely attributed to the AdaNP module, meaning that the backbone architecture has no impact. Therefore, when applied to larger models (e.g., transformers), the relative computational impact becomes negligible. Indeed, the complexity is primarily influenced by the number of constraints and the batch size, with the latter being particularly significant when inference is performed on CPU hardware (see Section 4.1 and Appendix A.2).
Besides constraint guarantee, ENFORCE also enhances the accuracy of the MLP (the $R^{2}$ increases from 0.994 to 0.999). At first sight, this may appear to be only a marginal improvement. However, advancing deep learning methods to achieve near-perfect performance remains an open challenge in essential fields such as healthcare and self-driving cars, where even slight inaccuracies at the decimal level can lead to catastrophic consequences [3, 4].
Interestingly, ENFORCE outperforms the MLP in data-scarce regions of the domain, which in this dataset correspond to the domain extremities (as shown in Fig. 3(b)). More generally, ENFORCE also performs better under data-scarcity conditions when the models are trained on uniformly sampled fractions of the dataset (Fig. 3(d)). This observation suggests that constrained learning may enhance data efficiency.
Table 1: Regression accuracy and constraint guarantee of ENFORCE when compared with a multilayer perceptron (MLP) and a soft-constrained NN (Soft). Results for $\lambda=0.5$, $\epsilon_{T}=10^{-4}$, and $\epsilon_{I}=10^{-6}$ are reported. The inference on 1000 samples is run on a CPU (11th Gen Intel(R) Core(TM) i7-1185G7 @ 3.00GHz, 4 Core(s)) to prove efficiency also in the absence of GPU hardware. The inference time is also reported for a batch of 1000 predictions.
MAPE [%] $R^{2}$ Avg Residual [%] Max Residual [%] Inference Time [s]
MLP 0.339 $\pm$ 0.083 0.994 $\pm$ 0.003 1.47 $\pm$ 0.33 17.13 $\pm$ 3.94 0.002 $\pm$ 0.000
Soft 0.944 $\pm$ 0.143 0.972 $\pm$ 0.002 1.55 $\pm$ 0.16 7.77 $\pm$ 0.40 0.002 $\pm$ 0.000
ENFORCE 0.060 $\pm$ 0.028 0.999 $\pm$ 0.000 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.008 $\pm$ 0.003
Effects of constrained learning
Notably, ENFORCE outperforms the MLP even before the projection steps, demonstrating superior performance using only the neural network backbone (Fig. 4(a), dashed-pink line). This can be attributed to the structure of the hard-constrained learning process, where the predictions are adjusted via projection to satisfy underlying constraints. Unlike soft-constrained methods, which only penalize constraint violations in the loss function, hard-constrained optimization incorporates projection-based adjustments that transform predictions to adhere strictly to the constraints. Consequently, after a few training steps, the model benefits from constrained learning, aligning its predictions more closely with valid regions of the solution space, resulting in improved predictions even before projection. Similar insights are also provided by Chen et al. (2021) [31]. Therefore, the constrained learning approach is likely to yield improved results even when AdaNP is omitted during inference to enhance computational efficiency. However, it should be noted that in this scenario, constraint satisfaction cannot be guaranteed.
(a) $y_{1}$
(b) $y_{2}$
(c) Constraint Satisfaction
(d) Data Scarcity
Figure 3: Prediction comparison between ENFORCE ($\lambda=0.5$, $\epsilon_{T}=10^{-4}$, $\epsilon_{I}=10^{-6}$) and a multilayer perceptron (MLP). ENFORCE enhances the overall accuracy and guarantees satisfaction for highly nonlinear constraints. ENFORCE consistently performs better than a standard MLP even when trained on uniformly sampled fractions of the training dataset.
Training dynamics
To understand the training dynamics of ENFORCE, we analyze the loss curves shown in Fig. 4(a), where the training data loss of ENFORCE is compared to the MLP. Being interested in the effect of hard-constrained learning and to ease the visualization, we do not report here the loss curve of the soft-constrained NN. In this case study, AdaNP contributes to the learning process from the very early iterations (Fig. 4(b), orange line), suggesting that the projection operations positively guide the optimization process. The combination of feasible predictions and minimization of projection displacement drives the learning process toward more optimal outcomes.
The modified loss function effectively guides the training process toward smaller projection displacements (Fig. 4(b), green dashed line). The displacement loss decreases consistently during training due to the influence of the penalty term in the loss function. Moreover, the depth of AdaNP progressively diminishes over training iterations down to $\sim$1 layer (Fig.4(b), orange line), due to (1) improved overall regression accuracy and (2) smaller projection displacement (i.e., a better linear approximation of the constraints). This adaptive behavior optimizes computational resources by adjusting to the required tolerance at each iteration. Furthermore, this decay in AdaNP depth is consistently observed across different training tolerance values, as illustrated in Fig.6(a).
(a) Loss NN vs. ENFORCE
(b) ENFORCE Losses and AdaNP
Figure 4: ENFORCE demonstrates significantly improved convergence, achieving lower loss compared to an unconstrained MLP. Enhanced training performances are reported for the backbone network of ENFORCE even before the action of AdaNP. This effect is enabled by the simultaneous minimization of the projection displacement (in green) and the action of the AdaNP module (in yellow). Note that we report average values across multiple runs, which explains why the depth of AdaNP appears as a step function with non-integer values.
Constrained learning hyperparameters (training)
We systematically analyze the influence of hyperparameters, such as the weighting factor $\lambda$ and the tolerance $\epsilon_{T}$, on the constrained learning process. Fig. 5 shows the influence of the hyperparameters on the accuracy of trained ENFORCE models evaluated on the test set.
The impact of the training tolerance $\epsilon_{T}$ on the model accuracy does not exhibit a clear trend, as its effect varies unpredictably with the weighting factors. Moreover, its influence is generally small compared to the variance of different training runs (Fig. 5). Intuitively, a smaller tolerance $\epsilon_{T}$ necessitates deeper AdaNP modules, resulting in higher computational costs due to the increased number of neural projection layers. This effect is visible in Fig. 6(a), where the depth of AdaNP during training is reported (i.e., number of projection layers). The average depth of AdaNP increases to accommodate stricter tolerances. For example, it expands from one to three layers as the tolerance $\epsilon_{T}$ is tightened from 1 to $10^{-5}$. Notably, in this case study, AdaNP operates with a minimum of one projection layer (i.e., when the tolerance is set to 1) and a maximum of 100. More importantly, the required depth tends to have a slower decay during training, if compared to using less strict tolerances (as visible in Fig. 6(a)). Larger tolerances result, on average, in shallow AdaNP layers (approximately one layer). This significantly reduces the training time associated with the projection operations. Along with the minor impact on overall accuracy, this observation suggests setting the training tolerance $\epsilon_{T}$ to less stringent requirements.
The regression accuracy is evidently affected by the choice of the loss weighting factor $\lambda$ (Fig. 5). Remarkably, unlike the challenging task of tuning weighting factors in soft-constrained methods [18], the constrained learning approach proposed here positively impacts accuracy regardless of the specific weighting factor chosen (as shown in Fig. 5, the accuracy of ENFORCE is consistently greater than that of a standard MLP). However, an inappropriate choice of this parameter can result in suboptimal outcomes (e.g., when $\lambda=2$ in Fig. 5). Therefore, careful tuning of this hyperparameter is warranted.
Constrained learning hyperparameters (inference)
During inference, ENFORCE dynamically adapts the depth of AdaNP to ensure an average tolerance below $\epsilon_{I}=10^{-6}$ in this case study. The required depth, however, also depends on the training parameters. Fig. 6(b) illustrates the number of NP layers needed to satisfy the constraint under varying $\lambda$ and $\epsilon_{T}$. The weighting factor is shown to reduce the required number of NP layers by half, with no additional cost during training. This phenomenon can be attributed to the fact that, in the absence of a penalty for projection displacement, the neural network is free to learn a function that, although potentially far from the actual one, results in projections that fall within the vicinity of the ground truth. This approach, however, necessitates multiple projections. In contrast, the penalty term drives the model to learn a function that is sufficiently close to the ground truth, thereby reducing the number of neural projections required. Increasing the value of $\epsilon_{T}$ impacts (positively) the depth of AdaNP at inference time when the displacement penalty factor is set to be small during training. This finding further supports the recommendation of employing shallow AdaNP modules during training, by relaxing the value of $\epsilon_{T}$.
We conclude that the weighting factor $\lambda$ plays an important role by balancing the contribution of the projection displacement error. On the other hand, enforcing strict satisfaction during training at arbitrary small tolerance $\epsilon_{T}$ does not necessarily improve the overall outcome.
Figure 5: Influence of constrained learning hyperparameters on the accuracy of ENFORCE on the test set (note that here we plot the inverse of the mean squared error (MSE)). The weighting factor $\lambda$ favors the learning process if appropriately tuned. Conversely, the training tolerance $\epsilon_{T}$ exhibits a small impact on performance, suggesting it can be set based on available resources. Overall, despite the choice of hyperparameters, ENFORCE is more accurate than an MLP with the same complexity, while also satisfying the underlying nonlinear constraint.
(a) Depth of AdaNP (number of neural projection layers) during training.
(b) Depth of AdaNP (number of neural projection layers) during inference.
Figure 6: Dynamic evolution of AdaNP during training and inference when different training hyperparameters are chosen. At training time, AdaNP is deeper as a smaller constraints tolerance $\epsilon_{T}$ is chosen.
6 Conclusions and outlook
We developed ENFORCE, a method designed to guarantee neural network prediction satisfies any nonlinear constraint $c(x,y)=0$ within a specified tolerance $\epsilon$, without relying on external solvers or incurring prohibitive computational overhead. The effectiveness of the method is demonstrated in an illustrative case study chosen for its interpretability and ease of visualization. We found that (1) ENFORCE ensures the satisfaction of nonlinear constraints to within a tolerance $\epsilon$ for both average and maximum residuals, (2) the overall accuracy improves from traditional unconstrained MLP, achieving near-perfect performance on the test set, and (3) the constrained learning hyperparameters influence the learning process toward optimal performance; however, our method shows superior performance compared the unconstrained counterpart for a wide range of hyperparamer settings. Additionally, we observed promising performance in data-scarcity conditions (Fig. 3(d)).
This work opens several research avenues toward developing robust NNs that strictly adhere to underlying system knowledge. First, the current method can be extended to handle piecewise-defined constraints and nonlinear inequality constraints. Additionally, the requirement for the constraint to be a $\mathcal{C}^{1}$ function could be relaxed by leveraging sub-gradients [55]. Finally, alternative (e.g., weighted) projection approaches could be explored to better account for the morphology and scaling of the constraints.
The method has the potential to address specific challenges or complement existing approaches, including those based on neural networks and other machine learning models for regression tasks (e.g., Gaussian processes and support vector machines). For instance, hard constraints can be combined with soft-constraint techniques, such as PINNs, to reliably solve PDEs [56]. Additionally, the method could enhance learning performance in partially annotated datasets by inferring missing information through available constraints. Finally, an interesting future direction could involve applying the AdaNP module to GenAI models, guiding the generation process toward domain-compliant samples, such as for synthetic data or image/video generation.
Limitations
In this paragraph. we highlight the main limitation of the proposed method. Firstly, the effectiveness of ENFORCE is highly dependent on the regression capabilities of the neural network backbone. When the model lacks sufficient complexity to achieve accurate predictions, ENFORCE provides limited benefit. Furthermore, the method becomes computationally intensive when applied to systems with a large number of constraints (e.g., more than 1000). This is due to the computational cost of each neural projection layer, which scales as $O(N^{3})$ because of the matrix inversion operation, where $N$ is the number of constraints. Further details are provided in Appendix A.2.
Acknowledgements
This research is supported by Shell Global Solutions International B.V., for which we express sincere gratitude.
Appendix A Appendix
A.1 Mathematical derivations
This section provides the key mathematical derivations underlying the proposed method. These derivations are intended to illustrate the theoretical foundations of the approach and to support some of the discussions presented in the main text.
Closed-form neural projection
We derive here the closed-form expression defining a neural projection layer in Def. 1 (Section 4.1). Given the linearized projection problem in Eq. 3, we can define the Lagrangian function as:
$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}\left(c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})\right)$ (6)
Then, a local optimum can be found by solving the KKT conditions (i.e., primal and dual feasibility):
$\nabla_{y}\mathcal{L}=(y-\hat{y})+\left.J^{T}_{y}c\right|_{x,\hat{y}}\lambda=0$ (7)
$c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
To simplify the notation, we define the linear system as:
$(y-\hat{y})+B^{T}\lambda=0$ (8)
$By-v=0$
where:
$B$ $=\left.J_{y}c\right|_{x,\hat{y}}\in\mathbb{R}^{N_{C}\times N_{O}}$ (9)
$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})\in\mathbb{R}^{N_{C}}$
Solving the system, we obtain a closed form for the neural projection layer:
$\tilde{y}=(I-B^{T}(BB^{T})^{-1}B)\hat{y}+B^{T}(BB^{T})^{-1}v$ (10)
Deviation from Newton’s method
To support the discussion raised in Section 4.1, we show how our method deviates from standard Newton’s method for solving nonlinear KKT conditions. Given a NLP:
$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}$ (11)
s.t. $\quad c(x,y)=0$
With associated Lagrangian function:
$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}c(x,y)$ (12)
The primal and dual feasibility can be derived as:
$\nabla_{y}\mathcal{L}=(y-\hat{y})+J^{T}_{y}c(x,y)\lambda=0$ (13)
$c(x,y)=0$
If we linearized the system according to the Newton’s iteration, we obtain:
$(y_{0}-\hat{y})+J^{T}_{y}c|_{y_{0},\lambda_{0}}\lambda_{0}+(y-y_{0})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{y_{0},\lambda_{0}}(y-y_{0})+\left.J^{T}_{y}c\right|_{y_{0},\lambda_{0}}(\lambda-\lambda_{0})=0$ (14)
$c(x,y_{0})+J_{y}c|_{y_{0},\lambda_{0}}(y-y_{0})=0$
Thus, assuming to center the linearization in the neural network prediction, i.e., $y_{0}=\hat{y}$, and choosing $\lambda_{0}=0$:
$(y-\hat{y})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})+\left.J^{T}_{y}c\right|_{\hat{y},0}\lambda=0$ (15)
$c(x,\hat{y})+\left.J_{y}c\right|_{\hat{y},0}(y-\hat{y})=0$
We can conclude that, essentially, our NP layer solves a similar linear system (Eq. 7) which does not comprise the term $\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})$, hence avoiding the computation of the Hessian tensor $\mathbf{H}_{y}c$.
A.2 Implementation details
Batch local projection
The computationally most expensive operation in the neural projection layer is the matrix inversion $(BB^{T})^{-1}$ (Def. 1 in Section 4.1), which has a complexity of $\mathcal{O}(N^{3})$. At inference time, $N=N_{C}$, since $B\in\mathbb{R}^{N_{C}\times N_{O}}$ and $BB^{T}\in\mathbb{R}^{N_{C}\times N_{C}}$. Hence, considering a number of constraint $N_{C}<10^{3}$, the matrix inversion is performed in $\sim 1$ million FLOPs, which is a negligible amount for most modern CPUs and GPUs. During training, assuming the use of batch gradient descent and defining the batch size ($BS$) as the number of data points processed in a single iteration, an equivalent number of matrix inversions must be performed. Thus, the computational cost for a single batch is apparently $\mathcal{O}(BS\times N_{C}^{3})$. To address this, we leverage parallel computing on GPUs by constructing a rank-3 tensor $\mathbf{B}\in\mathbb{R}^{BS\times N_{C}\times N_{O}}$ to hold $BS$ local matrices $B$. Similarly, a rank-2 tensor $V\in\mathbb{R}^{BS\times N_{C}}$ is built to store $BS$ local vectors $\mathbf{v}$. Modern deep learning libraries enable batch operations, such as matrix inversion, which reduce the effective complexity to $\mathcal{O}(N_{C}^{3})$ (i.e., scaling only with the number of constraints).
Given these conditions and the capabilities of current hardware, the neural projection operation remains computationally efficient even during training when the number of constraints is fewer than a few thousand. This facilitates new computational approaches for handling and enforcing nonlinear constraints in neural networks.

### v3

@@
-ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection
+ENFORCE: Nonlinear Constrained Learning with Adaptive-depth Neural Projection
 Abstract
-Ensuring neural networks adhere to domain-specific constraints is crucial for addressing safety and ethical concerns while also enhancing prediction accuracy.
+Ensuring neural networks adhere to domain-specific constraints is crucial for addressing safety and ethical concerns while also enhancing inference accuracy.
 Despite the nonlinear nature of most real-world tasks, existing methods are predominantly limited to affine or convex constraints.
-We introduce ENFORCE, a neural network architecture that guarantees predictions to satisfy nonlinear constraints exactly.
-ENFORCE is trained with standard unconstrained gradient-based optimizers (e.g., Adam) and leverages autodifferentiation and local neural projections to enforce any $\mathcal{C}^{1}$ constraint to arbitrary tolerance $\epsilon$.
-We build an adaptive-depth neural projection (AdaNP) module that dynamically adjusts its complexity to suit the specific problem and the required tolerance levels.
-ENFORCE guarantees satisfaction of equality constraints that are nonlinear in both inputs and outputs of the neural network with minimal (and adjustable) computational cost.
-Keywords Constrained learning  $\cdot$ Hard-constrained neural networks  $\cdot$ Trustworthy AI  $\cdot$ Physics-informed machine learning
+We introduce ENFORCE, a neural network architecture that uses an adaptive projection module (AdaNP) to enforce nonlinear equality constraints in the predictions.
+We prove that our projection mapping is 1-Lipschitz, making it well-suited for stable training.
+We evaluate ENFORCE on an illustrative regression task and for learning solutions to high-dimensional optimization problems in an unsupervised setting.
+The predictions of our new architecture satisfy $N_{C}$ equality constraints that are nonlinear in both the inputs and outputs of the neural network, while maintaining scalability with a tractable computational complexity of $\mathcal{O}(N_{C}^{3})$ at training and inference time.
 1 Introduction
-Neural networks (NNs) are the backbone of many recent advancements in artificial intelligence (AI), excelling in tasks such as natural language processing, image analysis, and scientific discovery due to their modularity, simplicity, and strong generalization capabilities.
+Neural networks (NNs) are the backbone of many recent advancements in artificial intelligence (AI), excelling in tasks such as natural language processing, image analysis, and scientific discovery due to their modularity, and strong generalization capabilities.
 However, their ability falls short when strict adherence to domain-specific constraints is required.
-Depending on the task, prior knowledge and rules about the system (e.g., from physics, safety, or ethics) are often available and are typically leveraged by humans in decision-making processes.
-In contrast, traditional data-driven methods such as NNs rely solely on data.
-Thus, trained NNs may be accurate on a training and test data set but still may not satisfy constraints exactly, leading to inconsistent predictions.
-This limitation not only generates substantial skepticism, hindering their adoption in real-world applications but can potentially result in erroneous or unfeasible outcomes when NNs are applied to decision-making tasks.
-Moreover, when prior domain knowledge is available and can be expressed as analytical expressions such as equations, ensuring that NNs adhere to this information is crucial to avoid the suboptimal utilization of valuable insights.
+Depending on the task, prior knowledge about the system (e.g., from physics, safety, or ethics) is often available and is typically leveraged by humans in decision-making processes.
+In contrast, data-driven methods such as NNs rely solely on data.
+Thus, trained NNs may be accurate on a training and test data set but still may not satisfy known constraints, leading to inconsistent predictions.
+This limitation not only generates substantial skepticism, hindering their adoption in real-world applications, but can also lead to erroneous or physically infeasible outcomes in decision-making processes.
+Moreover, when domain knowledge is available as analytical equations, ensuring that NNs adhere to this information is crucial to avoid a suboptimal utilization of expert insights and potentially reduce data demand [1].
 Enforcing strict constraints in NNs is a promising area of research for many fields.
-For example, in AI for Science, integrating first-principle laws ensures physically consistent models enabling insightful scientific discovery [1, 2].
-More broadly, constraining neural network predictions can have a transformative impact in domains where strict adherence to critical requirements is essential.
-This includes applications such as safety-critical systems (e.g., self-driving vehicles [3], healthcare [4]), fairness and bias mitigation [5, 6], and compliance with regulatory standards (e.g., in finance [7]).
-Additionally, with the rise of generative AI (GenAI), enforcing constraints on generation processes could mitigate potential risks and ensure trustworthy adherence to given criteria.
-Constraining the neural network output to adhere to strict rules is also beneficial to tackle traditional machine learning challenges, such as overfitting and underfitting in data-scarce regimes [8].
-It is also plausible that such constraints could enable some level of extrapolation in the presence of incomplete data or in holes of convex hulls [9, 10].
-Hence, the combination of enhanced accuracy and adherence to pre-existing domain knowledge could enable the safe use of NNs as surrogate models in computationally expensive applications [11].
-Enforcing constraints in NNs is not an easy task.
-Early approaches often relied on incorporating penalty terms into the loss function to reduce violations of physical laws [12].
+For example, in AI for Science, integrating first-principle laws ensures physically consistent models, enabling insightful scientific discovery [2, 3] or system modeling in engineering [4].
+A prominent area of application involves using NNs to safely accelerate computationally intensive tasks such as learning surrogate models [5] or solutions to (parametric) constrained optimization problems, either in a supervised or unsupervised manner [6, 7, 8, 9].
+More broadly, constraining neural network predictions can have a transformative impact in domains where strict adherence to critical requirements is essential, including safety-critical systems [10, 11], bias mitigation [12, 13], and compliance with regulatory standards [14].
+Additionally, with the rise of generative AI (GenAI), enforcing constraints on generation processes could mitigate risks by ensuring that generated data respects given criteria [15].
+Constraining the neural network output to adhere to strict rules is also beneficial to tackle traditional machine learning challenges, such as overfitting in data-scarce regimes [16].
+Enforcing constraints in NNs is not straightforward.
+The majority of existing approaches rely on incorporating penalty terms into the loss function to minimize constraint violations [17].
 Yet, these penalty-based methods offer no guarantees of constraint satisfaction (soft-constrained).
 In contrast, other methods aim to ensure strict adherence to analytical constraints by design (hard-constrained).
-For instance, one can use activation functions such as sigmoids or ReLU to bound outputs.
-To enforce analytical constraints beyond bounds, recent studies incorporate correction layers into NNs to project or complete predictions, ensuring they lie within the feasible region.
-Recent research has developed efficient methods to enforce analytical constraints defined by affine relationships between input and output variables or by convex regions (See Section 2).
-However, many real applications such as in science or sociology are inherently governed by nonlinear relationships between the involved variables [13, 14].
-Existing approaches for handling nonlinear constraints predominantly rely on external root-finding algorithms or constrained optimization solvers (Section 2.2).
-These methods introduce significant computational overhead and complicate model construction, thereby compromising the modularity and flexibility typically associated with NNs.
-In this paper, we propose ENFORCE, a neural network architecture that ensures predictions strictly satisfy nonlinear constraints representing domain knowledge or critical requirements.
-ENFORCE is trained using standard unconstrained optimization techniques and leverages an adaptive-depth projection module to enforce constraints by construction without relying on external solvers.
+For instance, one can use sigmoid functions to bound outputs.
+To enforce analytical constraints, recent studies incorporate correction layers into NNs to project or complete the predictions, ensuring they lie within a feasible region.
+For example, developed methods can enforce constraints defined by affine relationships between input and output variables or by convex regions [18, 19, 16].
+However, many applications, e.g., in science or sociology, are inherently governed by nonlinear constraints [20, 21].
+Existing approaches for handling nonlinear constraints predominantly rely on external root-finding or constrained optimization solvers [7, 22].
+These methods introduce significant computational overhead and complicate model development, thereby compromising the modularity and flexibility typically associated with NNs.
+In this paper, we propose ENFORCE, a neural network architecture that enforces predictions to satisfy nonlinear constraints.
+ENFORCE is trained using standard unconstrained optimization techniques and leverages an adaptive-depth neural projection (AdaNP) module to enforce constraints by construction without relying on external solvers.
 2 Related work
-This section reviews existing approaches to enforce constraints in NNs.
-According to the focus of this paper, a more in-depth analysis will be conducted on existing methods for enforcing algebraic equality constraints exactly.
+This section reviews different approaches to enforcing constraints in NNs, with a focus on existing hard-constrained methods.
 2.1 Soft-constrained neural networks
@@
 Soft constraints are incorporated as penalty terms appended to the loss function, penalizing residuals of algebraic [23, 24] or differential equations underlying the system [25].
-Physics-informed neural networks (PINNs) [12] represent a widely used framework designed to solve partial differential equations (PDEs) with NNs by employing soft constraints and collocation points.
+Physics-informed Neural Networks (PINNs) [17] represent a widely used framework designed to solve partial differential equations (PDEs) with deep learning by employing soft constraints and collocation points.
 Although the soft-constrained approach places no restrictions on the complexity of the constraints, it has the drawback of not guaranteeing strict adherence to them.
-Furthermore, increasing the complexity of the loss function – especially when the different terms vary in nature or scale – can degrade the optimization performance of the neural network, often resulting in suboptimal accuracy [18, 19].
+Furthermore, the complication of the loss landscape – especially when the different terms vary in nature or scale – can degrade the optimization performance of the neural network, often resulting in suboptimal accuracy [26, 27].
 2.2 Hard-constrained neural networks
-Hard-constrained NNs refer to methodological approaches ensuring that neural network predictions strictly adhere to analytical constraints.
-These constraints, explicitly encoded within the architecture, act as inductive biases, guiding the learning process toward compliance with domain knowledge or restrictions [20].
-Architectures such as convolutional neural networks (CNNs) [21] and graph neural networks (GNNs) [22, 23] encode inductive biases by guaranteeing invariance with respect to patterns and symmetries.
-Simple analytical constraints can be enforced using differentiable functions, such as sigmoids or ReLU for output bounding and softmax for simplex constraints (non-negativity and normalization).
-Recent literature includes significant contributions for enforcing analytical inequality constraints, such as convex polytopes and convex sets more generally [24, 25, 26, 27, 28].
-One can also constrain the neural network to guarantee specific functional characteristics, such as being 1-Lipschitz [29] or Lyapunov stable [30], nevertheless, this is not a core objective of this study.
-For a broad and recent review on hard-constrained NNs, the reader is also referred to [8].
+Hard-constrained neural networks refer to methodological approaches ensuring that neural network predictions adhere to given constraints by construction.
 Since this paper focuses on analytical equality constraints, the following literature review considers existing methods for this specific case.
@@
 Many methods for encoding hard equality constraints utilize projection techniques, which correct preliminary neural network predictions by appending a non-trainable layer to the output.
-Projections can be formulated as optimization problems (e.g., distance minimization) or derived from geometric principles.
-For example, in [31] NN predictions of physical systems governed by PDEs are projected to ensure solutions satisfy the finite difference discretization of the underlying linear PDEs.
-A more general approach is the KKT-hPINN [32], which enforces linear equality constraints of the form $Ax+By=b$, though its applicability is limited to affine equations.
+Projections can be formulated as optimization problems (i.e., distance minimization) or derived from geometric principles.
+For example, in [19] neural network predictions of physical systems governed by PDEs are projected to ensure solutions satisfy the finite difference discretization of the underlying linear PDEs.
+A more general approach is the KKT-hPINN [18], which enforces linear equality constraints in the inputs and outputs.
 Recently, HardNet [16] was introduced to enforce equality and inequality constraints affine in the output, without input restrictions, via a closed-form projection step.
-HardNet-Cvx [8] extends this approach to more general convex functions in the inputs, leveraging differentiable optimization layers supported by optimization solvers [33].
 Predict-and-complete
-Alternatively, the neural network can predict a partial set of output variables, $y_{P}\in\mathbb{R}^{N_{O}-N_{C}}$, and complete the prediction by solving the system of constraints based on this partial output.
+NNs can also predict a subset of output variables, $z\in\mathbb{R}^{N_{O}-N_{C}}$, and complete the prediction by solving the system of constraints based on this partial output (null-space methods).
 This approach ensures that the constraints are always satisfied.
@@
 introduced this concept to simulate physical systems such as climate modeling [28].
-However, when the constraints are not explicitly defined, solving the system requires a root-finding solver.
-Similar approaches have been proposed within the hybrid modeling community, particularly in the serial configuration, where a fully data-driven method is used to predict certain inputs to a mechanistic model [35].
+However, when the constraints are not available in explicit form, solving the system requires a root-finding solver.
+Similar approaches have been proposed within the hybrid modeling community, particularly in the serial configuration, where a fully data-driven method is used to predict unknown inputs to a mechanistic model [4].
 While studies like DC3 [7] have developed efficient backpropagation techniques, scenarios involving implicit nonlinear constraints can be computationally expensive to tackle with predict-and-complete methods.
-Moreover, the predict-and-complete approach can reduce regression accuracy by hindering the representation capabilities of NNs [34].
+Moreover, we rigorously show in Appendix B.6 that predict-and-complete approaches can suffer training instabilities if the constraints Jacobian is ill-conditioned [28].
 Constrained optimization
-One of the most popular fields of application of constrained learning is for solving (parametric) constrained optimization problems using NNs [36].
+To enforce analytical constraints, researchers leveraged constrained optimization to deploy specialized layers or directly train the neural network.
 OptNet [29] is an optimization layer developed to solve quadratic programs.
 Angrawal et al.
-[33] expands the methodology to convex programs.
+[30] expand the methodology to convex programs.
 They develop efficient differentiation techniques through such layers.
+Min et al.
+[16] leveraged such optimization layers to develop HardNet-Cvx, a neural network enforcing convex constraints.
 However, the forward pass always requires the solution of a constrained optimization problem.
-Recently, Mukherjee and Bhattacharyya [38] brute-forced the constrained learning paradigm by training a neural network using an NLP solver such as IPOPT [39] instead of standard unconstrained optimization algorithms.
-Also, NNs have been formulated as constrained optimization problems and solved using deterministic global solvers [11].
-However, these approaches pose severe limitations in terms of NNs and dataset size.
-Other methods
-Other methods have been proposed for constrained learning in NNs, mostly considering affine or convex regions.
-Many of them consider constraints only dependent on the input of the neural network [10, 27, 40, 41], others design strategies to include the dependence on both inputs and outputs [28, 42].
-Recently, contributions to enforce general logic and linear constraints have been proposed by the neuro-symbolic AI community, developing constraining layers using logic programming to improve multi-label classification and synthetic data generation [43, 44].
-However, to the best of our knowledge, there is no method designed to enforce any nonlinear equality constraints involving both the input and the output of the neural network by encoding them in the architecture and without relying on external solvers.
+Recently, Mukherjee and Bhattacharyya [22] approached the constrained learning paradigm by training a neural network using a constrained optimization solver such as IPOPT [31] instead of standard unconstrained optimization algorithms.
+However, these approaches pose severe limitations in terms of neural network and dataset size.
+Other methods have been proposed for constrained learning in NNs (cf.
+Appendix A).
+However, to the best of our knowledge, no existing method enforces nonlinear equality constraints involving both the input and output of a neural network by embedding them into the architecture while allowing training with unconstrained solvers.
 3 Preliminaries
 Problem statement
-Given a dataset $(x_{i},y_{i})_{i=1,...,N}$, we aim to train a neural network $f_{\theta}$ with parameters $\theta$ to approximate the underlying relationships while satisfying a set of known algebraic equality constraints $c(x,y)=0$.
-In general, $c$ can be any nonlinear function in the input $x$ and output $y$ of the neural network, incorporating domain knowledge or specifying critical requirements.
+Given a dataset $(x^{*}_{i},y^{*}_{i})_{i=1,...,N}$, without loss of generality we consider a neural network $f_{\theta}$ with parameters $\theta$ to approximate the underlying relationships while satisfying a set of known algebraic equality constraints $c(x,y)=0$.
+In general, $c$ can be a nonlinear function in the input $x$ and output $y$ of the neural network, incorporating domain knowledge or specifying critical requirements.
 Assumption 1.
-i) The constraints $c(x,y)=0$ are feasible, ii) $N_{C}<N_{O}$, where $N_{C}$ is the number of equality constraints and $N_{O}$ is the output dimensionality of the neural network, i.e., there are available degrees of freedom to learn.
-One way to enforce the neural network prediction $\hat{y}$ to satisfy the constraints is to project it onto the feasible hypersurface defined by $c(x,y)=0$.
-The projection can be defined as a mathematical program:
-$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{T}W\|y-\hat{y}\|$ (1)
-s.t. $\quad c(x,y)=0$
-[…]
+i) The constraints $c(x,y)=0$ are feasible and linearly independent, ii) $N_{C}<N_{O}$, where $N_{C}$ is the number of equality constraints and $N_{O}$ is the output dimensionality of the neural network, i.e., there are available degrees of freedom to learn.
+One way to enforce the neural network prediction $\hat{y}$ to satisfy the constraints is to project it onto the feasible hypersurface (manifold) defined by $c(x,y)=0$.
+The projection operation can be defined as an optimization problem:
+$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}(y-\hat{y})^{T}W(y-\hat{y})\quad\text{s.t.}\quad c(x,y)=0$ (1)
+If $W$ is the identity matrix, the prediction is corrected by an orthogonal projection onto the feasible region.
+This can be interpreted as finding the feasible solution that minimizes the Euclidean distance from the original prediction $\hat{y}$.
+A local solution to the nonlinear program in Eq.
+1 can be found by solving the first-order necessary optimality conditions, known as Karush–Kuhn–Tucker (KKT) conditions (cf.
+Chapter 12.3 in [33]).
+However, the latter is not necessarily straightforward as it may involve solving a system of nonlinear equations.
 Quadratic projection
-When $c(x,y)$ is an affine function in the neural network input and output (i.e., $c(x,y)=Ax+By-b=0$), then the problem results in a quadratic program (QP) and a closed-form analytical solution is available for the KKT conditions.
-Based on such observation, Chen et al.
-proposed KKT-hPINN by stacking a fixed non-trainable projection layer on top of a neural network, to guarantee the prediction adheres to given affine functions [32].
-However, the method is restricted to linear constraint as a closed-form is not available when $c$ is nonlinear.
-We note that a closed-form is still available when generalizing to any function in the input $x$, as the projection operation is still a QP.
-We observe that, assuming an affine dependence on $y$ such that $c=C(x)y-v(x)-b=0$, $C(x)$ and $v(x)$ play the roles of linear coefficients and translation, respectively, in a local sense.
-That is, given a input $x_{i}$ and a prediction $\hat{y}_{i}$, then any function of $x_{i}$ is a constant term with respect to the optimization problem in Eq.
-1, which becomes again a QP.
-Enforcing affine functions is not new and is achieved also through other techniques [8, 40].
-However, relaxing the assumption to allow for any nonlinear constraints of both the input and output of the neural network commonly results in decreased computational efficiency, as typically requires the use of constrained optimization techniques [38] or nonlinear solvers such as Newton’s methods [25].
+When $c(x,y)$ is an affine function in the neural network input and output, then the problem results in a quadratic program (QP) and a closed-form analytical solution is available for the KKT conditions [18].
+An extension to the closed-form is available when generalizing to any function in the input $x$, as the projection operation is still a QP.
+Consider an affine constraint on $y$ of the form $c=C(x)y-v(x)-b=0$, where $C(x)$ and $v(x)$ act as the linear coefficient matrix and translation vector, respectively.
+For a given input $x_{i}$ and prediction $\hat{y}_{i}$, any function of $x_{i}$ can be treated as constant with respect to the optimization problem in Eq.
+1, which thus reduces to a QP.
+Enforcing affine functions is not new and is also achieved through other techniques [16, 34].
+However, relaxing the assumption to allow for nonlinear constraints of both the input and output of the neural network commonly results in decreased computational efficiency and stability, as it typically requires the use of constrained optimization [22] or root-finding solvers such as Newton’s methods [7].
 4 Nonlinear constrained learning
 We present ENFORCE, a framework designed for general and efficient nonlinear constrained learning.
-The method employs a computationally cheap adaptive neural projection and has no restriction on the nonlinearity of $\mathcal{C}^{1}$ constraints involving both the input and output vectors of the neural network.
-At inference time, the model predictions satisfy the constraints to arbitrary tolerance $\epsilon$.
+The method employs a computationally cheap adaptive neural projection module and has no restriction on the nonlinearity of $\mathcal{C}^{1}$ constraints involving both the input and output of the neural network.
+We prove the neural projection to be a $1$-Lipschitz mapping, implying adversarial robustness and stable gradient flow dynamics when compared to state-of-the-art constrained learning methods such as predict-and-complete.
 4.1 AdaNP: Adaptive-depth neural projection
-We reformulate the NLP in Eq.
+We locally approximate the nonlinear program in Eq.
 1 and exploit the efficiency of quadratic projections to generalize the methodology to any nonlinear constraint.
@@
 where $J_{x}c$ and $J_{y}c$ are the Jacobian matrices with respect to the variable $x$ and $y$, respectively.
-Since the neural network input is fixed for a given sample, the approximation is exact in $x$, hence, $x=x_{0}$.
-Considering orthogonal projection for notation simplicity, the nonlinear optimization problem in Eq.
+Since the neural network input is fixed for a given sample, the linearization is exact in $x$, thus, $x=x_{0}$.
+Considering orthogonal projection, the nonlinear optimization problem in Eq.
 1 is locally approximated by a (linearly constrained) QP:
-$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{2}$ (3)
-s.t. $\quad c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
-Definition 1.
+$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}\|y-\hat{y}\|^{2}\quad\text{s.t.}\quad c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$ (3)
+Definition 1 (Projection operator $\mathcal{P}$).
 Given an input $x\in\mathbb{R}^{N_{I}}$ to a neural network $f_{\theta}$, its prediction $\hat{y}=f_{\theta}(x)\in\mathbb{R}^{N_{O}}$, and a set of constraints $c\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N{{}_{C}}})$, with $N_{C}<N_{O}$, we define an operator $\mathcal{P}$ such that $\tilde{y}=\mathcal{P}(\hat{y})$ is the solution to the linearized quadratic program in Eq.
 3, in the domain $\Omega$ where the constraints are defined.
-In particular, $\tilde{y}=B^{*}\hat{y}+v^{*}$, with $B^{*}=I-B^{T}(BB^{T})^{-1}B$ and $v^{*}=B^{T}(BB^{T})^{-1}v$, where $I\in\mathbb{R}^{N_{O}\times N_{O}}$ is the identity matrix, and:
-$B$ $=\left.J_{y}c\right|_{x,\hat{y}}$
-$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})$
-Since a closed-form expression for the operator $\mathcal{P}$ is derived (Appendix A.1), we can define a differentiable neural projection (NP) layer representing the operator $\mathcal{P}$.
-The forward and backward passes of an NP layer are computationally cheap (more details on implementation and computational cost are given in Appendix A.2).
+In particular, $\tilde{y}=B^{*}\hat{y}+v^{*}$, with $B^{*}=I-B^{T}(BB^{T})^{-1}B$ and $v^{*}=B^{T}(BB^{T})^{-1}v$, where $I\in\mathbb{R}^{N_{O}\times N_{O}}$ is the identity matrix, $B=\left.J_{y}c\right|_{x,\hat{y}}$, and $v=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})$.
+Given the closed-form expression of the operator $\mathcal{P}$ derived in Appendix B.1, we can define a differentiable neural projection (NP) layer representing the operator $\mathcal{P}$.
+The forward and backward passes of an NP layer are computationally cheap (more details on implementation and computational cost are given in Appendix D).
 However, the operator $\mathcal{P}$ projects the neural network prediction onto a linear approximation of the nonlinear constraints (i.e., the tangent hyperplane).
-The error that we introduce is proportional to the projection displacement $\ell_{p}=||\tilde{y}-\hat{y}||$.
-From this consideration, it follows that (1) the error is mitigated as the projection displacement is small, i.e., the neural network prediction is sufficiently accurate, and (2) a single NP layer may not ensure exact adherence to nonlinear constraints.
-Nevertheless, it is worth noting that a single NP layer guarantees strict satisfaction of equality constraints that are affine in $y$ and nonlinear in $x$, i.e., it efficiently enforces constraint classes considered in similar recent works [32, 8].
-To address the challenge of ensuring exact adherence to nonlinear constraints, as highlighted in the second consideration above, we propose AdaNP: an adaptive-depth neural projection composition to enforce nonlinear constraint satisfaction to arbitrary tolerance $\epsilon$.
-Definition 2.
+The error that we introduce is proportional to the projection displacement $e_{D}=||\tilde{y}-\hat{y}||$.
+From this consideration, it follows that (1) the error is mitigated as the projection displacement is small, i.e., the neural network prediction is sufficiently accurate, and (2) a single NP layer cannot ensure exact adherence to nonlinear constraints.
+It is worth noting that a single NP layer guarantees strict satisfaction of equality constraints that are affine in $y$ and nonlinear in $x$, i.e., it efficiently enforces constraint classes considered in similar recent works [18, 16].
+To address the challenge of satisfying nonlinear constraints, we propose AdaNP: an adaptive-depth neural projection composition that, under certain conditions, enforces nonlinear constraint satisfaction to arbitrary tolerance $\epsilon$.
+Definition 2 (AdaNP module).
 Given an operator $\mathcal{P}$ as defined in Def. 1, AdaNP is a composition of $n$ operators $\mathcal{P}$, such that:
 $\text{AdaNP}=\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n}$
-Then, given an arbitrarily small scalar $\epsilon$ and $n\in\mathbb{N}$ such that $|c(x,\tilde{y}_{n})|<\epsilon$, the feasible prediction $\tilde{y}_{n}$ is computed as:
-$\tilde{y}_{n}=(\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n})(\hat{y})$
+Proposition 1 (Convergence of AdaNP).
+Given an arbitrarily small scalar $\epsilon$, $n\in\mathbb{N}$ and assuming $\hat{y}$ in the positive reach (cf.
+Def.
+4, Appendix B.4) of the constraints manifold =$\mathcal{M}=\{x\in\mathbb{R}^{N_{I}},\>y\in\mathbb{R}^{N_{O}}:c(x,y)=0\}$, then $\tilde{y}_{n}$ is computed as:
+$\tilde{y}_{n}=(\mathcal{P}_{1}\circ\dots\circ\mathcal{P}_{n})(\hat{y}),$
+and converges to a feasible prediction such that $|c(x,\tilde{y}_{n})|<\epsilon$ with linear convergence rate under constraint smoothness conditions (cf. proof in Appendix B.3).
 AdaNP is a differentiable stack of $n$-NP layers that can be composed on every neural network backbone.
-The depth $n$ adjusts adaptively depending on the nonlinearities and the specified tolerance.
-Accurate NNs typically result in shallower AdaNP modules, since the projection displacement $\ell_{p}$ is related to the loss $\ell_{d}=||\hat{y}-y^{*}||$, where $y^{*}$ is the ground truth output.
+The depth $n$ adjusts adaptively during training and inference depending on the nonlinearities and the specified tolerance (cf.
+Algorithm 1 in Appendix D.1 for details about the adaptive behavior).
+Accurate NNs typically result in shallower AdaNP modules, since the linearization error $e_{D}$ is related to the distance $||\hat{y}-y^{*}||$ between the neural network prediction $\hat{y}$ and ground truth output $y^{*}$.
 This introduces a trade-off between the complexity of the backbone and the required depth of AdaNP to satisfy the specified tolerance criteria.
-It is worth mentioning that a similar concept of iterative neural projections, although employed for a different application and objective, is also utilized in [45].
-Analogy with Sequential Quadratic Programming
-AdaNP can also be seen as an iterative method that recursively improves the solution of a linearized nonlinear program.
-Here, we notice the similarity to sequential quadratic programming (SQP) techniques.
-Specifically, AdaNP is a simple case of SQP method for which the objective function is naturally quadratic while the nonlinear constraints are linearized (in contrast to full SQP, in which the objective function is quadratically approximated).
-This observation could potentially allow to analyze the convergence of the method using SQP theory [46, 47].
 Deviation from Newton’s method
-While the KKT conditions for an NLP (Eq.
-1) can be more generally solved using Newton’s methods, our method circumvents the computational overhead associated with calculating the Hessian matrix of the constraints (Appendix A.1).
+While the KKT conditions for a nonlinear program (Eq.
+1) can be more generally solved using Newton’s methods, our method circumvents the computational overhead associated with calculating the Hessian matrix of the constraints (cf.
+Appendix B.2) at the cost of reduced convergence rate (i.e., full Newton’s method converges quadratically).
+4.2 Conditioning analysis
+The set of constraints $c(x,y)=0$ describes an infinite-wide feasible region (i.e., a hypersurface) where the constraints are defined.
+Hence, one could ask whether the projections are unique as well as whether the projection mapping is stable and well-conditioned, and thus suitable for guiding the learning process.
+We provide theoretical conditions for the uniqueness and no-worse property of the projected prediction, along with evidence of regularity and robustness, and show that other state-of-the-art constrained learning approaches do not guarantee these properties.
+Proposition 2 (Uniqueness of the projection).
+We consider the submanifold defined by the linearly independent constraints in the output space =$\mathcal{M}=\{y\in\mathbb{R}^{N_{O}}:c(y)=0\}$, where $c$ are real-valued, continuously differentiable functions with Lipschitz-continuous gradients (hence, the manifold has positive reach [35]).
+If $\hat{y}\in\operatorname{reach}(\mathcal{M})$, then there exists a unique orthogonal projection of $\hat{y}$ onto $\mathcal{M}$ (cf.
+Appendix B.4 and [35]).
+The reach depends on the geometry of the constraint set (e.g., the reach is infinite for linear constraints and approaches zero near sharp corners or singularities).
+Therefore, for the projection to be well-defined, $\hat{y}$ must lie within a sufficiently small tubular neighborhood of $\mathcal{M}$.
+Proposition 3 (No-worse property).
+When a unique projection $\tilde{y}$ exists and the manifold $\mathcal{M}$ has positive reach, then the projection mapping satisfies the no-worse property: for any prediction $\hat{y}$ and ground truth $y^{*}$, $||y^{*}-\tilde{y}||\leq||y^{*}-\hat{y}||$, that is, the projected prediction is always a no-worse prediction than the original one (cf.
+Appendix B.5).
+Regularity and robustness of appended layers (e.g., projection operation or null-space completion) can influence the stability of the neural network training.
+We prove (cf.
+Appendix B.6) the projection operation to be a non-expansive mapping in the neighborhood of the constraint manifold (i.e., its Lipschitz constant $L\leq 1$).
+This ensures adversarial robustness of the neural projection layer during the forward pass and stable gradient flow dynamics.
+Theorem 1 (Non-expansiveness of the projection operator).
+Given the nonlinear program:
+$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}\quad\text{s.t.}\quad c(y)=0,$ (4)
+with $\hat{y}\in\mathbb{R}^{N_{O}}$, $\tilde{y}\in\mathbb{R}^{N_{O}}$, and $c(y):\mathbb{R}^{N_{O}}\rightarrow\mathbb{R}^{N_{C}}$ smooth, continuous constraints with a full-rank Jacobian $J_{y}c$, the projection operator $\mathcal{P}:\tilde{y}=\mathcal{P}(\hat{y})$ solving the nonlinear program (Eq.
+4) is a non-expansive mapping in the neighborhood of the constraints manifold:
+$\forall\hat{y}_{i},\hat{y}_{j}\in\mathbb{R}^{N_{O}},$ $\operatorname{dist}(\hat{y}_{i},\mathcal{M})<\operatorname{reach}(\mathcal{M})\land\operatorname{dist}(\hat{y}_{j},\mathcal{M})<\operatorname{reach}(\mathcal{M})$ (5)
+$\Rightarrow\ \left\|\mathcal{P}(\hat{y}_{i})-\mathcal{P}(\hat{y}_{j})\right\|\leq\left\|\hat{y}_{i}-\hat{y}_{j}\right\|$
+Lemma 1 (Gradient flow dynamics).
+The non-expansive property of the projection mapping ensures the stability of the gradient flow dynamics.
+During backpropagation, the pre-projection gradient is multiplied by a factor $J_{\hat{y}}\mathcal{P}^{T}$, which is the bounded Lipschitz constant.
+If $\nabla_{\tilde{y}}\ell$ is the post-projection gradient, then we can bound the pre-projection gradient:
+$\nabla_{\hat{y}}\ell=J_{\hat{y}}\mathcal{P}^{T}\nabla_{\tilde{y}}\ell\>\Rightarrow\>||\nabla_{\hat{y}}\ell||\leq||J_{\hat{y}}\mathcal{P}^{T}||\,||\nabla_{\tilde{y}}\ell||$ (6)
+Thus, no exploding gradients can originate from the projection step, as the scale of pre-projection and post-projection gradients is comparable.
+Conversely, alternative constrained learning paradigms do not guarantee bounds on the operated mapping.
+For example, predict-and-complete approaches using Newton-based root-finding methods (e.g., DC3 [7]) can exhibit non-Lipschitz behavior when the constraint Jacobian is ill-conditioned (cf.
+Appendix B.6), potentially leading to unstable training and slower convergence.
 4.3 Architecture
 The architecture of ENFORCE (Fig.
-1) is composed of (1) a backbone neural network, that can be of any kind and complexity, and (2) an AdaNP module.
+1) is composed of (1) a neural network (without loss of generality) as backbone, which can be of any kind and complexity, and (2) an AdaNP module.
 The depth of AdaNP depends on the backbone performance and specified tolerance.
-Indeed, AdaNP can be purposely tuned to increase training efficiency (Section 4.4).
+Indeed, the tolerance of AdaNP can be tuned to increase training efficiency (cf.
+Appendix D.1).
 A single NP layer is composed of two steps: (1) automatic differentiation and (2) local neural projection.
@@
 Exact Jacobian computation
-To compute the Jacobian of the constraint system, we leverage automatic differentiation available in most deep learning libraries [48, 49, 50].
+To compute the Jacobian of the constraint system, when not available analytically, we leverage automatic differentiation available in most deep learning libraries [36, 37, 38].
 Computing the local Jacobian $J_{y}c|_{x,\hat{y}}$ is computationally inexpensive, as it requires propagating derivatives only through the constraints and does not involve the neural network backbone.
@@
 Thus, the projection is locally defined in the neighborhood of ($x_{i},\hat{y}_{i}$).
-We parallelize the computation of local neural projections by building a rank-3 tensor $\mathbf{B}^{*}$ and a rank-2 tensor $\mathbf{v}^{*}$ (Appendix A.2).
-Thus, we reduce the cost of an NP layer from $\mathcal{O}(BS\times N_{C}^{3})$ to $\mathcal{O}(N_{C}^{3})$, allowing effective training with stochastic gradient descent techniques with virtually no limitation on the batch dimension ($BS$).
-For modern hardware, handling output dimensionalities of $N_{C}<10^{3}$ leads to minimal computational cost.
-4.3 Implications of orthogonal projections
-The set of constraints $c(x,y)=0$ describes an infinite-wide feasible region (i.e., a hypersurface) where the constraints are defined.
-Hence, one could ask whether the projections are unique and thus suitable for guiding the learning process.
-To this end, theoretical insights can be derived by analyzing the case of orthogonal projections, i.e., $W=I$ (Eq.
-1).
-(a) Affine constraint
-(b) Nonlinear constraint
-Figure 2: Orthogonal projection on the feasible region when $y\in\mathbb{R}^{2}$.
-Orthogonal projections on affine constraints are unique and always lead to enhanced predictions.
-When the constraint is nonlinear, multiple orthogonal projections are possible, not necessarily improving the prediction.
-Affine constraints
-Considering constraints that are affine in $y$, the feasible space is a hyperplane defined in $\mathbb{R}^{N_{O}}$.
-Under those circumstances, it can be proven that there exists a unique orthogonal projection defining $\tilde{y}$ from a point $\hat{y}\in\mathbb{R}^{N_{O}}$ such that $c(\tilde{y})=0$.
-More interestingly, if $y^{*}$ is the ground truth, then $\forall\hat{y},y^{*},\,||y^{*}-\tilde{y}||<||y^{*}-\hat{y}||$, that is, the projected prediction is always a better prediction than the original one (Fig.
-2(a)).
-In such a condition, constrained learning effectively guides the learning process toward the ground truth.
-In particular, the loss is expected to decrease compared to the unconstrained case from the very start of the training process.
-Nonlinear constraints
-The same considerations do not hold for nonlinear hypersurfaces.
-Firstly, there could be multiple points on the hypersurface where the normal vectors point toward the prediction $\hat{y}$.
-Secondly, the closest orthogonal projection can potentially represent a worse prediction, since the distance $||y^{*}-\tilde{y}||$ is not ensured to be shorter than $||y^{*}-\hat{y}||$ (see Fig.
-2(b)).
-Then, the projection step could potentially conflict with the training step, preventing the training process itself from converging.
-Therefore, we can not conclude that orthogonal projections onto a nonlinear feasible region always guide the learning process efficiently.
+We parallelize the computation of local neural projections by building a rank-3 tensor $\mathbf{B}$ and a rank-2 tensor $\mathbf{v}$ (Appendix D.3).
+Thus, we reduce the apparent complexity of an NP layer from $\mathcal{O}(BS\times N_{C}^{3})$ to $\mathcal{O}(N_{C}^{3})$, allowing effective training with stochastic gradient descent techniques.
+On modern hardware, handling up to $N_{C}<10^{3}$ constraints results in a computational cost that remains practical, particularly when using Cholesky decomposition for matrix inversion [39].
+Moreover, the complexity of this method is equivalent to other state-of-the-art methods such as DC3 [7].
+Additional insights into memory requirements are provided in Appendix D.4.
 4.4 Training ENFORCE
-We train ENFORCE using standard unconstrained gradient descent methods (i.e., Adam).
-We develop and use a constrained learning methodology using AdaNP to guide the neural network training to convergence, supported by the theoretical implications of orthogonal projections described above (Section 4.3).
-Linearized constraints and training strategy
-ENFORCE does not project the predictions directly on the nonlinear feasible region but on a locally linearized counterpart.
-Thus, we can assume the existence of a unique closest orthogonal projection.
-However, the quality of the latter, i.e., whether the projection really improves the prediction, is largely influenced by the constraint geometry curvature.
-Moreover, the initialization of the neural network plays a crucial role in constrained learning using ENFORCE.
-Overall, it is difficult to prove a priori whether constrained learning can be beneficial during the initial phase of training.
-To overcome this limitation, we develop a strategy to ensure effective constrained learning.
-Instead of building a static ENFORCE model, we activate AdaNP only when its application enhances prediction accuracy (i.e., the loss decreases).
-This strategy can be interpreted as an unconstrained pre-training; however, it is automatically incorporated in ENFORCE.
-We have demonstrated that this approach accelerates the training process and leads to better optima compared to fully unconstrained training.
+We train ENFORCE using standard unconstrained gradient descent methods (i.e., Adam [32]).
+We develop and use a constrained learning methodology using AdaNP to guide the neural network training to convergence, supported by the theoretical implications of orthogonal projections described in Section 4.2.
 Loss function
-We further control the constrained learning process by minimizing the projection displacement, i.e., the distance between the preliminary NN prediction $\hat{y}$ and its projection $\tilde{y}=\text{AdaNP}(\hat{y})$.
-This aims to (1) ensure minor linearization error ($\epsilon_{L}\sim\Delta y$) and (2) prevent the neural network from learning alternative functions whose projections fall within the neighborhood of the desired functions.
-The latter is suggested to reduce reliance on AdaNP for accurate predictions, thereby lowering the computational cost during inference (i.e., by decreasing the depth of AdaNP).
-We control the projection displacement by applying soft constraints, adding an extra penalty term to the loss function:
-$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\tilde{y}_{i}||^{2}+\frac{\lambda}{N}\sum_{i=1}^{N}||\hat{y}_{i}-\tilde{y}_{i}||^{2}$ (4)
-Here, the first term represents the conventional mean squared error between the projected prediction $\tilde{\mathbf{y}}_{i}$ and the ground truth $\mathbf{y}_{i}$, while the second term quantifies the projection displacement, targeted at controlling and minimizing its magnitude, by acting as a penalty.
-The two terms are balanced by a scalar weight $\lambda$.
+The loss function used throughout this study takes the following general form:
+$\ell=\ell_{T}+\ell_{D}+\ell_{C}=\ell_{T}+\frac{\lambda_{D}}{N}\sum_{i=1}^{N}||\hat{y}_{i}-\tilde{y}_{i}||^{2}+\frac{\lambda_{C}}{N}\sum_{i=1}^{N}||c(x_{i},\tilde{y}_{i})||,$ (7)
+where the first term, $\ell_{T}$, is a task-specific loss function selected based on the target model.
+The second and third terms are regularization penalties that respectively minimize the projection displacement, $||\hat{y}_{i}-\tilde{y}_{i}||$, and the constraint residual $||c(x_{i},\tilde{y}_{i})||$.
+The relative contributions of these terms are controlled by the scalar weights $\lambda_{D}$ and $\lambda_{C}$.
+Minimizing the projection displacement aims to (1) ensure minor linearization error ($\epsilon_{L}\sim\Delta y$) and (2) prevent the neural network from learning alternative functions whose projections onto the constraints fall within the neighborhood of the desired functions.
+Also, this additional loss term is suggested to reduce reliance on AdaNP, thereby lowering the computational cost during inference (i.e., by decreasing the depth of AdaNP).
+Adaptive training strategy
+We propose a strategy to facilitate constrained learning during the early stages of training, guided by the theoretical insights presented in Section 4.2.
+In the initial training phases, the preliminary prediction $\hat{y}$ may be inaccurate and lie far from the constraint manifold $\mathcal{M}$.
+Under such conditions, projecting onto a locally linearized approximation of the constraints can introduce substantial errors in the prediction.
+To mitigate this issue in practice, inspired by trust-region methods [33], we activate AdaNP only when the projection operation leads to an improvement in the prediction accuracy (e.g., quantified by a decrease in some loss measure $m_{\ell}$).
+This often leads to an unconstrained pre-training phase, followed by the activation of the AdaNP module.
+In other words, this serves as a heuristic to ensure that the prediction $\hat{y}$ lies sufficiently close to the constraint manifold $\mathcal{M}$.
+Details about the underlying algorithm and loss measures used in this study are reported in Appendix D.2.
 5 Experiments and discussion
-We demonstrate our proposed method on an illustrative two-dimensional example. We aim to fit the oscillating function $y:\mathbb{R}\rightarrow\mathbb{R}^{2}$:
+We evaluate the proposed method on two tasks: (i) fitting a vector-valued function with implicitly and nonlinearly related components, and (ii) learning solutions to nonlinear parametric optimization problems.
+While the function fitting case study is kept low-dimensional for illustrative purposes, we benchmark ENFORCE on high-dimensional optimization problems following the state-of-the-art experimental routine in DC3 [7].
+All experiments were conducted using an NVIDIA A100 Tensor Core GPU 80 GB, while the nonlinear programming solver (Section 5.2) runs on a CPU (11th Gen Intel(R) Core(TM) i7, 4 Core(s), 8 Logical Processor(s)).
+5.1 Illustrative function fitting
+We aim to fit the illustrative oscillating function $y:\mathbb{R}\rightarrow\mathbb{R}^{2}$:
 $y_{1}$ $=2\text{sin}(fx)$ (8)
 $y_{2}$ $=-\text{sin}^{2}(fx)-x^{2}\text{,}$
-where $x\in\mathbb{R}$ is the unidimensional independent variable (input) and the scalar $f$ is the frequency.
+where $x\in\mathbb{R}$ is the unidimensional independent variable (input) and the scalar $f=5$ is the frequency.
 Notably, the system is implicitly linked by a nonlinear constraint $c(x,y_{1},y_{2})=(0.5y_{1})^{2}+x^{2}+y_{2}$, involving both input and output variables.
-We train an ENFORCE model consisting of a 64-neuron 1-hidden layer fully connected ReLU neural network as a backbone and an AdaNP module to force the predictions to satisfy the constraint.
-We experiment with different constraint tolerances ($\epsilon_{T}$) and loss weighting factors ($\lambda$) during training to heuristically study their influence.
-At inference, the depth of AdaNP is adapted to guarantee a specified tolerance $\epsilon$, e.g., absolute average or maximum residual across a batch of prediction.
-In this case study, we set the average absolute residual to be below $\epsilon_{I}=10^{-6}$ at inference time.
-To verify the interpolation capabilities, we sample 100 training data points from a uniform distribution in $x=[-2,2]$ and 1000 test points in the same domain.
+We train an ENFORCE model consisting of a 64-neuron 1-hidden-layer fully connected ReLU neural network as a backbone and an AdaNP module to force the predictions to satisfy the constraint.
+The supervised task loss is the mean squared error ($\ell_{T}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\tilde{y}_{i}||^{2}$), while $\lambda_{C}$ is set to zero (i.e., the constraint is addressed exclusively by AdaNP and no soft constraint term is used).
+At inference time, the depth of AdaNP is adaptively adjusted to enforce the constraint within a specified tolerance $\epsilon$, such as the absolute average or maximum residual across a batch of predictions.
+In this case study, we set the maximum absolute residual to be below $\epsilon_{I}=10^{-6}$ at inference time.
+To verify the regression capabilities, we sample 100 training data points from a uniform distribution in $x=[-2,2]$ and 100,000 test points in the same domain.
+We compare the method with a multilayer perceptron (MLP) and a soft-constrained neural network sharing the same architecture.
+We report here the outcomes of the best-performing soft-constrained network, with a constraint loss weight $\lambda_{C}=1$.
+We train all the models for 50,000 epochs, using Adam optimizer [32] and a learning rate of $10^{-3}$.
 Every run is repeated 5 times using different initialization seeds.
-Accuracy and constraint guarantee
+Table 1: Regression accuracy and constraint guarantee of ENFORCE on 100,000 test samples when compared with a multilayer perceptron (MLP) and a soft-constrained neural network (Soft).
+Results for $\lambda_{D}=0.5$, $\epsilon_{T}=10^{-4}$, and $\epsilon_{I}=10^{-6}$ are reported.
+We report the inference time for a batch of 1,000 samples.
+(Note that $\text{MAPE}=\frac{100\%}{N}\sum_{i=1}^{N}\left|\frac{y^{*}_{i}-\tilde{y}_{i}}{y^{*}_{i}}\right|$).
+Method MAPE [%] $R^{2}$ Mean eq. [%] Max eq. [%] Inf. time [s]
+MLP 0.339 $\pm$ 0.083 0.994 $\pm$ 0.003 1.47 $\pm$ 0.33 17.13 $\pm$ 3.94 0.002 $\pm$ 0.000
+Soft 0.944 $\pm$ 0.143 0.972 $\pm$ 0.002 1.55 $\pm$ 0.16 7.77 $\pm$ 0.40 0.002 $\pm$ 0.000
+ENFORCE 0.060 $\pm$ 0.028 0.999 $\pm$ 0.000 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.008 $\pm$ 0.003
+We report the main results in Table 1.
+Notably, ENFORCE guarantees constraint satisfaction (below the required tolerance) with minimal computational overhead (+6 ms) while also improving predictive accuracy over standard MLPs, addressing the core challenge of balancing expressivity and feasibility in constrained learning.
+Extensive discussion about performance, constrained learning dynamics, and hyperparameters is reported in Appendix C.
+Notably, we observe that, after constrained training, ENFORCE outperforms the MLP even before projection (i.e., using just the backbone at inference).
+5.2 Constrained optimization problem
+A relevant field in neural network research involves learning approximate solutions to large-scale constrained optimization problems.
+In this context, NNs serve as fast and inexpensive alternatives to traditional solvers, which are often computationally intensive.
+Following the benchmarking protocol introduced by Donti et al.
+[7], we first evaluate ENFORCE on learning the solution of nonconvex optimization problems with linear equality constraints.
+We then extend the evaluation by replacing the linear equality constraints with nonlinear ones.
+We compare against alternative methods for learning (or solving) constrained optimization problems, such as MLP, soft constrained MLP, the state-of-the-art DC3 [7], and the deterministic nonlinear programming solver IPOPT [31].
+For every model, we use an equivalent fully connected ReLU neural network backbone consisting of 2 hidden layers with 200 neurons.
+Training is performed with a batch size of 200, a learning rate of $10^{-4}$, and until model convergence (3,500 epochs for DC3; 1,000 for ENFORCE and the other baselines).
+Every run is repeated 5 times.
+Like Donti et al., we run the test inference on a single batch of 833 samples, and for the inference time of the optimizer, we assume full parallelization on 833 CPUs (cf.
+[7] for more details).
+5.2.1 Nonconvex problem with linear constraints
+We consider the same class of nonconvex optimization problems as in Donti et al. [7], with the focus on equality constraints:
+$\min_{y\in\mathbb{R}^{N_{O}}}f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y},\quad\text{s.t.}\quad c^{T}y=x,$ (9)
+where $Q\in\mathbb{R}^{N_{O}\times N_{O}}\succeq 0$, $p\in\mathbb{R}^{N_{O}}$, and $c\in\mathbb{R}^{N_{C}\times N_{O}}$ are randomly sampled constant parameters, while $x\in\mathbb{R}^{N_{C}}$ (with $N_{C}$ = $N_{I}$) is the variable parameter across problem instances.
+$Q$ is a diagonal matrix chosen to be positive semi-definite and $x$ is uniformly sampled in the interval $[-5,5]$.
+We aim to learn the optimal $y$ given an instance of $x$ in an unsupervised fashion.
+Rather than using a dataset of solved optimization instances, we minimize the objective in the unsupervised task loss $\ell_{T}=f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y}$.
+In Table 2, we report the results on the constrained nonconvex task for 200 variables and 150 linear equality constraints.
+Given the linear nature of the constraints, ENFORCE consistently guarantees the feasibility for all the test samples.
+ENFORCE learns a solution that is only $0.47\%$ suboptimal relative to the objective value obtained by IPOPT, while obtaining a 25×25$\times$25 × acceleration.
+ENFORCE shows faster training convergence when compared to DC3 (cf.
+Table 4, Appendix C) and an optimal objective gain of $40\%$.
+This improvement can be attributed to the stability of the projection mapping, in contrast to the null-space completion method.
+As expected, unconstrained and soft-constrained methods do not guarantee feasibility and may yield predicted optima with lower objective values than those computed by constraint-respecting solvers.
+However, in constrained optimization, such infeasible solutions are inadmissible, regardless of their objective value.
+In Appendix C (Table 4), we show how the performance of the methods scales with varying numbers of variables and constraints, including training time.
+Table 2: Results on a batch of 833 instances of nonconvex optimization problems with linear equality constraints involving 200 variables and 150 equality constraints.
+ENFORCE guarantees the feasibility of the solutions, is 25×25$\times$25 × faster than IPOPT, although assuming full parallelization, and learns a $40\%$ better optimum than the state-of-the-art DC3 method.
+Baseline deep learning-based and soft-constrained methods show significant constraint violations and suboptimal predictions.
+Method Obj. value Max eq. Mean eq. Inf. time [s]
+IPOPT [31] $-10.64\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.379\pm 0.060$
+MLP $-52.99\pm 0.01$ $45.38\pm 0.56$ $9.14\pm 0.02$ $0.001\pm 0.001$
+Soft ($\lambda_{c}=5$) $1.28\pm 0.32$ $1.45\pm 0.43$ $0.08\pm 0.00$ $0.001\pm 0.000$
+Soft ($\lambda_{c}=1$) $-8.18\pm 0.18$ $1.47\pm 0.41$ $0.09\pm 0.00$ $0.001\pm 0.001$
+Soft ($\lambda_{c}=0.1$) $-13.55\pm 0.02$ $2.17\pm 0.12$ $0.35\pm 0.00$ $0.001\pm 0.001$
+DC3 [7] $-6.27\pm 0.07$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.004\pm 0.000$
+ENFORCE $-10.59\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.016\pm 0.002$
+5.2.2 Nonconvex problem with nonlinear constraints
+We extend the evaluation to the following class of problems involving nonlinear constraints in both inputs and outputs, and perform high-dimensional analysis:
+$\min_{y\in\mathbb{R}^{N_{O}}}f_{\text{obj}}(y)=\frac{1}{2}y^{T}Qy+p^{T}\sin{y},\quad\text{s.t.}\quad y^{T}Ay+c^{T}y+d=x^{3}.$ (10)
+Here, $A\in\mathbb{R}^{N_{C}\times N_{O}\times N_{O}}$ denotes a tensor holding $N_{C}$ randomly generated symmetric matrices, while the remaining parameters follow the same sampling procedure as previously described.
+The varying parameter $x$ is uniformly drawn from the range $[-5,5]$, and the dataset dimensionality and split remain unchanged.
+We consider a problem with 200 variables and 150 nonlinear constraints.
+The results are reported in Table 3.
+ENFORCE successfully predicts optimal solutions satisfying the set of nonlinear constraints across the whole test set.
+ENFORCE consistently achieves a 25× speedup in inference compared to the nonlinear programming solver, while maintaining the optimal objective within $6\%$.
+Traditional deep learning and soft constraint methods perform poorly when faced with nonlinear constraints, resulting in significant infeasibility or failure to approximate an optimal solution.
+Note that a comparison with DC3 is not included, as the original implementation does not support large-scale problem benchmarking involving nonlinear constraints.
+Once again, a scalability analysis across multiple problem dimensions is reported in Appendix C (Table 5).
+Table 3: Performance comparison on nonconvex problems with nonlinear equality constraints involving 200 variables and 150 constraints.
+ENFORCE predicts feasible and near-optimal solutions across the entire test set, achieving a 25×25$\times$25 × speedup over IPOPT.
+Traditional deep learning and soft-constraint methods exhibit significant constraint violations and suboptimal performance.
+Method Obj. value Max eq. Mean eq. Inf. time [s]
+IPOPT [31] $-29.45\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $3.40\pm 1.40$
+MLP $-53.07\pm 0.00$ $497.38\pm 4.64$ $118.39\pm 0.07$ $0.002\pm 0.001$
+Soft ($\lambda_{c}=5$) $(3.02\pm 0.04)\times 10^{5}$ $79.23\pm 3.81$ $16.72\pm 0.07$ $0.001\pm 0.001$
+Soft ($\lambda_{c}=1$) $(6.03\pm 0.04)\times 10^{4}$ $79.30\pm 3.70$ $16.68\pm 0.06$ $0.001\pm 0.000$
+Soft ($\lambda_{c}=0.1$) $(6.11\pm 0.01)\times 10^{3}$ $78.35\pm 1.95$ $16.63\pm 0.01$ $0.002\pm 0.001$
+ENFORCE $-27.77\pm 0.02$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.14\pm 0.08$
+6 Conclusions
+We propose ENFORCE, a method to ensure that neural network predictions satisfy a set of $\mathcal{C}^{1}$ nonlinear constraints $c(x,y)=0$ within a specified tolerance $\epsilon$, without relying on external solvers or incurring significant computational overhead.
+We prove the stability of the proposed projection mapping during training and provide theoretical insights into its convergence properties and applicability.
+The effectiveness of the method is demonstrated heuristically on (i) an illustrative function fitting case study selected for its interpretability and ease of visualization, and (ii) large-scale nonconvex optimization problems with nonlinear constraints.
+Our findings show that (1) ENFORCE consistently achieves constraint feasibility up to specified tolerance in the case studies, (2) task performance improves (up to 40%) when using ENFORCE over baseline methods in both function fitting and learning-based optimization settings, and (3) ENFORCE accelerates (×$\times 25$) computationally intensive tasks such as constrained optimization.
+Acknowledgments and Disclosure of Funding
+This research is supported by Shell Global Solutions International B.V., for which we express sincere gratitude.
+Appendix A Additional related works
+Hard-constrained neural networks belong to the broader category of techniques incorporating inductive biases in deep learning models.
+Inductive biases guide the learning process toward compliance with domain knowledge or restrictions [40].
+Architectures such as convolutional neural networks (CNNs) [41] and graph neural networks (GNNs) [42, 43] encode inductive biases by guaranteeing invariance with respect to patterns and symmetries.
+Simple analytical constraints can be enforced using differentiable functions, such as sigmoids or ReLU for output bounding and softmax for simplex constraints.
+Recent literature includes significant contributions for enforcing analytical inequality constraints, such as convex polytopes and convex sets more generally [44, 7, 45, 46, 47].
+One can also constrain the neural network to guarantee specific functional characteristics, such as Lipschitz continuity [48] or Lyapunov stability [49].
+Other methods have been proposed for constrained learning in NNs with analytical and logical expressions, mostly considering affine or convex regions [50, 51].
+Many of them consider constraints only dependent on the input of the NN [52, 46, 34, 53], others design strategies to include the dependence on both inputs and outputs [47, 5].
+Recently, contributions to enforce general logic and linear constraints have been proposed by the neuro-symbolic AI community, developing loss terms or constraining layers using logic programming [54, 55, 56].
+For a broad and recent review on hard-constrained NNs, the reader is also referred to [16].
+Appendix B Mathematical derivations
+This section provides the key mathematical derivations, theorems, and proofs underlying the proposed method.
+These derivations are intended to illustrate the theoretical foundations of the approach and to support some of the discussions presented in the main text.
+B.1 Closed-form neural projection
+We derive here the closed-form expression defining a neural projection layer in Def. 1 (Section 4.1). Given the linearized projection problem in Eq. 3, we can define the Lagrangian function as:
+$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}\left(c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})\right)$ (11)
+Then, a local optimum can be found by solving the KKT conditions (i.e., primal and dual feasibility):
+$\nabla_{y}\mathcal{L}=(y-\hat{y})+\left.J^{T}_{y}c\right|_{x,\hat{y}}\lambda=0$ (12)
+$c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
+To simplify the notation, we define the linear system as:
+$(1)$ $(y-\hat{y})+B^{T}\lambda=0$ (13)
+$(2)$ $By-v=0$
+where:
+$B$ $=\left.J_{y}c\right|_{x,\hat{y}}\in\mathbb{R}^{N_{C}\times N_{O}}$ (14)
+$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})\in\mathbb{R}^{N_{C}}$
+Solving the system, we obtain a closed form for the neural projection layer:
+$\tilde{y}=(I-B^{T}(BB^{T})^{-1}B)\hat{y}+B^{T}(BB^{T})^{-1}v$ (15)
+B.2 Deviation from Newton’s method
+To support the discussion raised in Section 4.1, we show how our method deviates from standard Newton’s method for solving nonlinear KKT conditions. Given a nonlinear program:
+$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}$ (16)
+s.t. $\quad c(x,y)=0$
+With associated Lagrangian function:
+$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}c(x,y)$ (17)
+The primal and dual feasibility can be derived as:
+$\nabla_{y}\mathcal{L}=(y-\hat{y})+J^{T}_{y}c(x,y)\lambda=0$ (18)
+$c(x,y)=0$
+Linearizing the system according to Newton’s iteration at $y=y_{0}$ results in:
+$(y_{0}-\hat{y})+J^{T}_{y}c|_{y_{0},\lambda_{0}}\lambda_{0}+(y-y_{0})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{y_{0},\lambda_{0}}(y-y_{0})+\left.J^{T}_{y}c\right|_{y_{0},\lambda_{0}}(\lambda-\lambda_{0})=0$ (19)
+$c(x,y_{0})+J_{y}c|_{y_{0},\lambda_{0}}(y-y_{0})=0$
+Thus, assuming to center the linearization in the neural network prediction, i.e., $y_{0}=\hat{y}$, and choosing $\lambda_{0}=0$:
+$(y-\hat{y})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})+\left.J^{T}_{y}c\right|_{\hat{y},0}\lambda=0$ (20)
+$c(x,\hat{y})+\left.J_{y}c\right|_{\hat{y},0}(y-\hat{y})=0$
+We can conclude that, essentially, our NP layer solves a similar linear system (Eq.
+12) which does not comprise the term $\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})$, hence avoiding the computation of the Hessian tensor $\mathbf{H}_{y}c$.
+Here, we note some similarity with Gauss-Newton methods used to solve least square problems [33].
+B.3 Local convergence rate
+AdaNP can also be seen as an iterative method that recursively improves the solution of a linearized nonlinear program.
+Here, we notice the similarity to sequential quadratic programming (SQP) techniques.
+Specifically, AdaNP is a simple case of SQP method for which the objective function is naturally quadratic while the nonlinear constraints are linearized (in contrast to full SQP, in which the objective function is quadratically approximated).
+This observation allows to analyze the convergence rate of the method starting from SQP theory [33, 57, 58].
+The projection operator $\mathcal{P}$ solves an SQP subproblem of the form:
+$\min_{y}$ $\frac{1}{2}(y-\hat{y})^{T}\bar{H}(y-\hat{y})$ (21)
+s.t. $\quad J_{y}c\,(y-\hat{y})+c(x,\hat{y})=0,$
+which approximates the original nonlinear program:
+$\min_{y}$ $\frac{1}{2}(y-\hat{y})^{T}I(y-\hat{y})$ (22)
+s.t. $\quad c(x,y)=0$
+The Hessian $\bar{H}$, as observed in the deviation from Newton’s method B.2, does not include the second-order derivatives of the constraints that would appear in the full Lagrangian Hessian.
+The resulting method is often called Gauss-Newton SQP step, since the way the constraints derivatives are dropped reminds of the Gauss-Newton method for nonlinear least squares [33].
+Supported by SQP theory [33], conditions for local convergence can be derived.
+We assume $y^{*}$ to be a local solution to the original nonlinear program (Eq.1) at which the following conditions hold [33]:
+  1. H1
+    The objective function and the constraints are twice differentiable in a neighborhood of $y^{*}$ with Lipschitz continuous second derivatives.
+  2. H2
+    The linear independence constraint qualification (LICQ) holds at $y^{*}$. Then, the KKT conditions are satisfied for a vector of Lagrangian multipliers $\lambda^{*}$.
+  3. H3
+    The second-order sufficient conditions (SOSC) hold at ($y^{*}$, $\lambda^{*}$).
+The KKT conditions for the original nonlinear program are defined as:
+$F(z)=\begin{bmatrix}\nabla_{y}\mathcal{L}(y,\lambda)\\ c(y)\end{bmatrix},\quad\text{with}\quad z=\begin{bmatrix}y\\ \lambda\end{bmatrix},$ (23)
+and are satisfied by a vector $z^{*}=\begin{bmatrix}y^{*}&\lambda^{*}\end{bmatrix}^{T}$.
+We define the Jacobian of the KKT conditions of the original nonlinear program (Eq. 22) in a neighborhood of the local solution as:
+$J^{(k)}=\begin{bmatrix}\nabla_{yy}^{2}\mathcal{L}^{(k)}&J_{c}^{T}\\ J_{c}&0\end{bmatrix}$ (24)
+We assume that LICQ and SOQC hold also in the neighborhood of $z^{*}$ (H2 and H3), hence the Jacobian at iteration $k$, $J^{(k)}$, is non-singular and thus invertible.
+The deviation of the projection operator $\mathcal{P}$ from the complete SQP step can be expressed through a matrix $E$ holding the second-order derivatives of the constraints:
+$E=\begin{bmatrix}\sum_{i}\lambda^{(k)}_{i}\nabla^{2}c_{i}(y^{(k)})&0\\ 0&0\end{bmatrix},$ (25)
+such that $J^{(k)}=\bar{J}+E$, with $\bar{J}$ being the Jacobian of the KKT conditions associated with the problem in Eq 21.
+At iteration $k$, we define the residual $r^{(k)}=F(z^{(k)})$, the error $e^{(k)}=z^{(k)}-z^{*}$ and solve for the Newton’s step $s^{(k)}$:
+[…]
+$z^{(k+1)}=z^{(k)}+s^{(k)}$
+$e^{(k+1)}=e^{(k)}+s^{(k)}$
+Since $F(z)$ is twice continuously differentiable, using Taylor expansion:
+$F(z^{(k)})=J^{(k)}e^{(k)}+r^{(k)},\quad\quad\text{with}\quad r^{(k)}=\mathcal{O}(||e^{(k)}||^{2})$ (27)
+Thus, a QP step can be expressed as:
+$\bar{J}s^{(k)}=-r^{(k)}=-F(z^{(k)})=-J^{(k)}e^{(k)}-r^{(k)}$ (28)
+From the definition of the Jacobian $\bar{J}$ and given that $J^{(k)}$ is invertible:
+$(J^{(k)}-E)s^{(k)}=-J^{(k)}e^{(k)}-r^{(k)}$ (29)
+$((J^{(k)})^{-1}J^{(k)}-(J^{(k)})^{-1}E)s^{(k)}=-(J^{(k)})^{-1}J^{(k)}e^{(k)}-(J^{(k)})^{-1}r^{(k)}$ (30)
+$(I-M)s^{(k)}=-e^{(k)}-(J^{(k)})^{-1}r^{(k)},$ (31)
+$s^{(k)}=-(I-M)^{-1}(e^{(k)}+(J^{(k)})^{-1}r^{(k)}),$ (32)
+with $M=(J^{(k)})^{-1}E$.
+Thus, in the neighborhood of the solution, the error at iteration $k+1$ can be expressed as:
+$e^{(k+1)}=z^{(k+1)}-z^{*}=e^{(k)}+s^{(k)}=e^{(k)}-(I-M)^{-1}(e^{(k)}+(J^{(k)})^{-1}r^{(k)})$ (33)
+Rearranging:
+$e^{(k+1)}$ $=(I-(I-M)^{-1})e^{(k)}-(I-M)^{-1}(J^{(k)})^{-1}r^{(k)}$ (34)
+$=(I-M)^{-1}((I-M)-I)e^{(k)}-(I-M)^{-1}(J^{(k)})^{-1}r^{(k)}$
+$=-(I-M)^{-1}Me^{(k)}-(I-M)^{-1}(J^{(k)})^{-1}r^{(k)}$
+Banach’s lemma then gives:
+$||(I-M)^{-1}||\leq\frac{1}{1-||M||}=\frac{1}{1-\rho}=C_{0}$ (35)
+Then we can estimate the error:
+$||e^{(k+1)}||\leq C_{0}(||M||\>||e^{(k)}||+||(J^{(k)})^{-1}||\>||r^{(k)}||)$ (36)
+Since $r^{(k)}=O(||e^{(k)}||^{2})$, $\exists\>C_{1}>0:||r^{(k)}||\leq C_{1}||e^{(k)}||^{2}$, then:
+$||e^{(k+1)}||\leq C_{0}||M||\>||e^{(k)}||+C_{0}||(J^{(k)})^{-1}||\>C_{1}||e^{(k)}||^{2})$ (37)
+We can conclude that, in the neighborhood of the solution:
+  • If $M=0$, the linear term vanishes and yields quadratic convergence, i.e., when the constraints are affine and thus the second order derivative of the constraints vanish ($\nabla^{2}c_{i}=0$).
+  • If $M\neq 0$ but $||M||<1$, it is guaranteed strictly linear convergence with rate $||M||$, plus a higher-order correction.
+  • If $||M||\geq 1$, the Gauss–Newton step alone may not converge. Second-order corrections (or using the full Lagrangian Hessian) are then required.
+Among state-of-the-art methods for constrained learning, we recognize that Newton-based completion approaches exhibit a quadratic convergence rate.
+However, as will be shown below (Appendix B.6), they can suffer from training instabilities.
+B.4 Uniqueness of the projection
+Given a set of constraints $c(y)=0$, sufficiently smooth and with full rank matrix $J_{y}c$ when $c(y)=0$, we define the $(N_{O}-N_{C})$-dimensional submanifold =$\mathcal{M}=\{y\in\mathbb{R}^{N_{O}}:c(y)=0\}$ in the ambient-space manifold $\mathcal{N}\in\mathbb{R}^{N_{O}}$.
+Then, we can prove that in the neighborhood of the manifold, the orthogonal projection is unique despite the nonlinearity of the constraints.
+Definition 3 (Tubular neighbourhood).
+Let $\mathcal{M}$ be a smooth embedded submanifold of a smooth manifold $\mathcal{N}$ and let $\nu(\mathcal{M})\to\mathcal{M}$ be its normal bundle.
+A tubular neighbourhood of $\mathcal{M}$ in $\mathcal{N}$ is an open set $U\subseteq\mathcal{N}$ for which there exists an open neighbourhood $\mathcal{V}$ of the zero section in $\nu(\mathcal{M})$ and a diffeomorphism $\Phi$:
+$\Phi:\mathcal{V}\;\longrightarrow\;U$
+that restricts to the identity on the zero section. In other words, $U$ is obtained by smoothly “thickening” $\mathcal{M}$ along its normal directions.
+Definition 4 (Reach of a manifold [35, Def. 4.1]).
+The reach of a closed subset $\mathcal{M}$ (in particular, an embedded submanifold), $\operatorname{reach}(\mathcal{M})$, is the largest radius $\rho>0$ for which $\forall\hat{y}$ such that $\operatorname{dist}(\hat{y},\mathcal{M})$ there is a unique nearest point $\tilde{y}\in\mathcal{M}$.
+Formally:
+$U_{\rho}:=\bigl{\{}y\in\mathbb{R}^{N_{O}}:\operatorname{dist}(y,\mathcal{M})<\rho\bigr{\}}.$
+For every $\rho>0$ define the nearest–point (metric) projection:
+$\mathcal{P}_{\rho}:U_{\rho}\,\longrightarrow\,\mathcal{M},\qquad\mathcal{P}_{\rho}(y)\;=\;\operatorname*{arg\,min}_{y\in\mathcal{M}}\|y-\hat{y}\|.$
+Since $\mathcal{M}$ is closed, the minimum exists, but it may fail to be unique.
+The reach of $\mathcal{M}$ is the supremum $\rho$ such that there is a single minimum to the projection above:
+[…]
+For instance, $\operatorname{reach}(\mathcal{M})=\infty$ exactly when the projection $\mathcal{P}_{\rho}$ is single-valued for every $\rho>0$, e.g. when $\mathcal{M}$ is an affine subspace.
+Theorem 2 (Uniqueness of the projection).
+Given a prediction $\hat{y}\in\mathbb{R}^{N_{O}}$ and a smooth constraints $c(y):\mathbb{R}^{N_{O}}\rightarrow\mathbb{R}^{N}_{C}$ with full rank Jacobian $J_{y}c$ when the constraints are satisfied, defining a submanifold =$\mathcal{M}=\{y\in\mathbb{R}^{N_{O}}:c(y)=0\}$, if $\operatorname{dist}(\hat{y},\mathcal{M})<\operatorname{reach}(\mathcal{M})$, then the minimizer $\tilde{y}$ defined as:
+$\tilde{y}=\operatorname*{arg\,min}_{y\in\mathcal{M}}\frac{1}{2}\|y-\hat{y}\|.$
+exists and is unique.
+Proof: The proof follows easily from Definition 3 and Definition 4.
+B.5 No-Worse prediction
+We prove that, when a unique projection exists on a constraint manifold with positive reach, if $y^{*}$ is the ground-truth, then the projected prediction $\tilde{y}$ is always a better prediction than the original prediction $\hat{y}$.
+Proposition 4 (No-worse property).
+Let $\mathcal{M}\subset\mathbb{R}^{N_{O}}$ be a smooth, embedded submanifold defined by equality constraints $c(y)=0$, with $c:\mathbb{R}^{N_{O}}\rightarrow\mathbb{R}^{N_{C}}$ smooth and $J_{y}c$ full rank and Lipschitzian on $\mathcal{M}$ (hence, $\mathcal{M}$ has positive reach [35]).
+Let $\operatorname{reach}(\mathcal{M})>0$.
+Suppose:
+  • $y^{*}\in\mathcal{M}$ is the (unknown) ground truth,
+  • $\hat{y}\in\mathbb{R}^{N_{O}}$ is a model prediction such that $\operatorname{dist}(\hat{y},\mathcal{M})<\operatorname{reach}(\mathcal{M})$,
+  • $\tilde{y}:=\mathcal{P}(\hat{y})$ is the orthogonal projection of $\hat{y}$ onto $\mathcal{M}$.
+Then:
+$\|\tilde{y}-y^{*}\|\leq\|\hat{y}-y^{*}\|,$
+with equality if and only if $\hat{y}\in\mathcal{M}$.
+That is, projecting onto the constraint manifold never increases the Euclidean error with respect to the ground truth, and strictly reduces it when the constraints are violated.
+Proof: Since $\hat{y}\in U_{\rho}:=\{y\in\mathbb{R}^{N_{O}}:\operatorname{dist}(y,\mathcal{M})<\operatorname{reach}(\mathcal{M})\}$, the orthogonal projection $\mathcal{P}(\hat{y})$ is well-defined and unique.
+Denote $\tilde{y}:=\mathcal{P}(\hat{y})$ and recall that $\tilde{y}$ is the closest point to $\hat{y}$ on $\mathcal{M}$.
+By the normality property of projections onto manifolds with positive reach (cf. [35, Def. 4.1]), we have:
+$\langle\hat{y}-\tilde{y},y^{*}-\tilde{y}\rangle\geq 0\quad\text{for all }y^{*}\in\mathcal{M}.$ (38)
+Expanding the squared distance yields:
+$\|\hat{y}-y^{*}\|^{2}$ $=\|\tilde{y}-y^{*}+\hat{y}-\tilde{y}\|^{2}$
+$=\|\tilde{y}-y^{*}\|^{2}+\|\hat{y}-\tilde{y}\|^{2}+2\langle\hat{y}-\tilde{y},\tilde{y}-y^{*}\rangle.$
+Using the projection property $\langle\hat{y}-\tilde{y},y^{*}-\tilde{y}\rangle\geq 0$, we get:
+$\|\hat{y}-y^{*}\|^{2}\geq\|\tilde{y}-y^{*}\|^{2}+\|\hat{y}-\tilde{y}\|^{2},$ (39)
+and hence:
+$\|\tilde{y}-y^{*}\|^{2}\leq\|\tilde{y}-y^{*}\|^{2}-\|\hat{y}-\tilde{y}\|^{2}$ (40)
+$\|\tilde{y}-y^{*}\|^{2}\leq\|\tilde{y}-y^{*}\|^{2}$
+$\|\tilde{y}-y^{*}\|\leq\|\hat{y}-y^{*}\|$
+Equality holds if and only if $\|\hat{y}-\tilde{y}\|=0$, i.e., $\hat{y}=\tilde{y}\in\mathcal{M}$.
+B.6 Conditioning of the projection operation
+The training stability and robustness can potentially be influenced by layers appended on top of a neural network (e.g., projection operation or null-space completion).
+We demonstrate that, in the neighborhood of the constraints manifold, the projection operation is a non-expansive mapping in the Banach sense, i.e., its Lipschitz constant $L\leq 1$.
+This implies stability and adversarial robustness during the forward pass and well-conditioned gradient-flow dynamics during the backward pass.
+On the other hand, null-space completion methods (i.e., in DC3 and other methods [7, 28]) are characterized by a Lipschitz constant dependent on the Jacobian of the constraints.
+This can lead to training instabilities, vanishing or exploding gradients, whenever the Jacobian is ill-conditioned.
+Theorem 3 (Non-expansiveness of the projection operator).
+Given a nonlinear program:
+$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}$ (41)
+s.t. $\quad c(y)=0,$
+with $\hat{y}\in\mathbb{R}^{N_{O}}$, $\tilde{y}\in\mathbb{R}^{N_{O}}$, and $c(y):\mathbb{R}^{N_{O}}\rightarrow\mathbb{R}^{N_{C}}$ are smooth, continuous constraints with a full-rank Jacobian $J_{y}c$.
+The projection operator $\mathcal{P}:\tilde{y}=\mathcal{P}(\hat{y})$ is a non-expansive mapping in the neighborhood of the constraints manifold:
+$\forall\hat{y}_{i},\hat{y}_{j}\in\mathbb{R}^{N_{O}},$ $\operatorname{dist}(\hat{y}_{i},\mathcal{M})<\operatorname{reach}(\mathcal{M})\land\operatorname{dist}(\hat{y}_{j},\mathcal{M})<\operatorname{reach}(\mathcal{M})$ (42)
+$\Rightarrow\ \left\|\mathcal{P}(\hat{y}_{i})-\mathcal{P}(\hat{y}_{j})\right\|\leq\left\|\hat{y}_{i}-\hat{y}_{j}\right\|$
+Proof: The proof begins by formally proving that $\mathcal{P}$ is an orthogonal projection operator, by studying its dynamics with respect to $\hat{y}$.
+The Lagrangian of the problem 41 can be expressed as:
+$\mathcal{L}(y,\lambda)=\frac{1}{2}||y-\hat{y}||^{2}+\lambda^{T}c(y)$ (43)
+At the optimal point $\tilde{y}=\mathcal{P}(\hat{y})$:
+$\begin{cases}y-\hat{y}+J_{y}c^{T}\lambda=0\\ c(y)=0\end{cases}$ (44)
+We now consider $y$ and $\lambda$ as implicit functions of $\hat{y}$:
+$y=y(\hat{y}),\quad\lambda=\lambda(\hat{y})\quad\Rightarrow F(y(\hat{y}),\lambda(\hat{y}),\hat{y})=\begin{cases}y(\hat{y})-\hat{y}+J_{y(\hat{y})}c^{T}\lambda(\hat{y})=0\\ c(y(\hat{y}))=0\end{cases}$ (45)
+The total derivative of $F$ can be derived as:
+$\frac{d}{d\hat{y}}(F)=\left.\frac{dF}{dy}\right|_{\hat{y},\lambda}\frac{dy}{d\hat{y}}+\left.\frac{\partial F}{\partial\lambda}\right|_{\hat{y},y}\frac{d\lambda}{d\hat{y}}+\left.\frac{dF}{d\hat{y}}\right|_{y,\lambda},$ (46)
+where:
+$\left.\frac{dF}{dy}\right|_{\hat{y},\lambda}=\begin{bmatrix}I+\frac{d}{dy}\left(J_{y(\hat{y})}c^{T}\lambda(\hat{y})\right)\\ J_{y}c\end{bmatrix}=\begin{bmatrix}I+H\\ J_{y}c\end{bmatrix}$ (47)
+$\left.\frac{\partial F}{\partial\lambda}\right|_{\hat{y},y}=\begin{bmatrix}J_{y}c^{T}\\ 0\end{bmatrix}$ (48)
+$\left.\frac{dF}{d\hat{y}}\right|_{y,\lambda}=\begin{bmatrix}-I\\ 0\end{bmatrix}$ (49)
+Rearranging in (block) matrix form:
+$\begin{bmatrix}\left.\frac{dF}{dy}\right|_{\hat{y},\lambda}&\left.\frac{\partial F}{\partial\lambda}\right|_{\hat{y},y}\end{bmatrix}\begin{bmatrix}\frac{dy}{d\hat{y}}\\ \frac{d\lambda}{d\hat{y}}\end{bmatrix}=-\begin{bmatrix}\left.\frac{dF}{d\hat{y}}\right|_{y,\lambda}\end{bmatrix}$ (50)
+$\begin{bmatrix}I+H&J_{y}c^{T}\\ J_{y}c&0\end{bmatrix}\begin{bmatrix}\frac{dy}{d\hat{y}}\\ \frac{d\lambda}{d\hat{y}}\end{bmatrix}=\begin{bmatrix}I\\ 0\end{bmatrix}$ (51)
+We simplify the notation to solve the linear system by defining:
+$Z$ $:=\frac{dy}{d\hat{y}}\in\mathbb{R}^{N_{O}\times N_{O}}$ (52)
+$L$ $:=\frac{d\lambda}{d\hat{y}}\in\mathbb{R}^{N_{C}\times N_{O}}$
+$J$ $:=J_{y}c\in\mathbb{R}^{N_{C}\times N_{O}}$
+Note that the row space of $J$ is the normal space to the constraints, while its kernel (column space) is the tangent space.
+Then:
+$\begin{cases}(I+H)Z+J^{T}L=I\\ JZ=0\end{cases}$ (53)
+The second equation means that every column of $Z$ lives in $\operatorname{ker}(J)$, i.e., the tangent space.
+Then, since $\hat{y}$ is in the reach of the constraint manifold, let $P:=I-J^{T}(JJ^{T})^{-1}J$ be the unique projector onto the subspace $\operatorname{ker}(J)$.
+$P$ satisfies the following properties:
+  • $P^{2}=P\Rightarrow$ idempotent
+  • $P^{T}=P\Rightarrow$ symmetric
+  • $JP=0\Rightarrow\operatorname{Im}(P)=\operatorname{ker}(J)$ ($\operatorname{Im}(\cdot)$ is the image and $\operatorname{ker}(\cdot)$ is the kernel)
+Then, any solution to the system in Eq. 53 must satisfy $Z=PZ$.
+Sub-proof: $Jz=0\Rightarrow$ each column $z^{(k)}$ lies in $\operatorname{ker}(J)$. Then:
+[…]
+Now, we can guess the simplest representative of $Z$ of the form of $PZ$, such as $Z=P$. Then, from the first equation and using the definition of the projector matrix:
+$(I+H)P+J^{T}L=I=P+J^{T}(JJ^{T})^{-1}J=P+Q,$ (55)
+where $Q=J^{T}(JJ^{T})^{-1}J$ is the normal component of the identity matrix. Developing the algebra:
+$IP+HP+J^{T}L=P+Q$ (56)
+$J^{T}L=Q-HP$ (57)
+Since $J$ has full rank, there exists a left-inverse (Moore-Penrose pseudoinverse) $J^{+}$ such that $J^{+}J=I$ and $J^{+}=(JJ^{T})^{-1}J$. Hence, solving for $L$ and using the definition of $Q$:
+$L$ $=J^{+}(Q-HP)$ (58)
+$=(JJ^{T})^{-1}JQ-(JJ^{T})^{-1}JHP$
+$=(JJ^{T})^{-1}JJ^{T}(JJ^{T})^{-1}J-(JJ^{T})^{-1}JHP$
+$=(JJ^{T})^{-1}J(I-HP)$
+We can conclude:
+$Z$ $=\frac{dy}{d\hat{y}}=I-J^{T}(JJ^{T})^{-1}J$ (59)
+$L$ $=\frac{d\lambda}{d\hat{y}}=(JJ^{T})^{-1}J(I-HP)$
+Considering now the projection operator $\mathcal{P}$ such that $y=\mathcal{P}(\hat{y})$, we found that the Jacobian of such operator is the orthogonal projector:
+$\frac{dy}{d\hat{y}}=J_{\hat{y}}\mathcal{P}=I-J^{T}(JJ^{T})^{-1}J$ (60)
+Hence, $J_{\hat{y}}\mathcal{P}$ is idempotent, symmetric, and does not depend on the second-order derivatives of the constraints in $H$.
+We can easily prove that all the eigenvalues of a symmetric and idempotent matrix are either 0 or 1.
+Thus, the spectral norm of the Jacobian is 1:
+$||J_{\hat{y}}\mathcal{P}||=\max\{|\lambda_{p}|\}=1,$ (61)
+with $\lambda_{i}$ being the $p$-th eigenvalue (to be not confused with the Lagrangian multipliers).
+Being the Jacobian bounded to 1, the projection operator $\mathcal{P}$ is a non-expansive mapping in $\mathbb{R^{N_{O}}}$.
+In other words, the mapping operated by $\mathcal{P}$ is $1$-Lipschitz:
+$\left\|\mathcal{P}(\hat{y}_{i})-\mathcal{P}(\hat{y}_{j})\right\|\leq\left\|\hat{y}_{i}-\hat{y}_{j}\right\|,$ (62)
+which concludes the proof of the theorem.
+Practical implications
+The bounding of the projection operator leads to important practical consequences in training a neural network, such as:
+  • Adversarial robustness (or forward sensitivity): Small input perturbation cannot be magnified by the projection operation, which follows directly from Eq. 62.
+  • Stability of gradient flow dynamics (or backward sensitivity): During backpropagation, the pre-projection gradient is multiplied by a factor $J_{\hat{y}}\mathcal{P}^{T}=J_{\hat{y}}\mathcal{P}$, which is bounded.
+    $\nabla_{\hat{y}}\ell=J_{\hat{y}}\mathcal{P}^{T}\nabla_{\tilde{y}}\ell\quad\text{with}\quad||\nabla_{\hat{y}}\ell||\leq||J_{\hat{y}}\mathcal{P}^{T}||\,||\nabla_{\tilde{y}}\ell||,$ (63)
+    with $\ell$ being the loss term.
+Hence, the gradient flow is well conditioned (no exploding gradients can originate from the projection step, the scale of pre-projection and post-projection gradients is comparable), which leads to smoother and faster training convergence.
+Comparison with null-space methods (predict-and-complete)
+In null-space methods, a neural network $f_{\theta}$ outputs a vector $z\in\mathbb{R}^{N_{O}-N_{C}}:z=f_{\theta}(x)$.
+Then, a mapping $\varphi_{x}$ uses the constraints $c(x,y)$ to complete the output vector with $\varphi_{x}(z)\in\mathbb{R}^{N_{C}}$ such that $y=\begin{bmatrix}z&\varphi_{x}(z)\end{bmatrix}^{T}$.
+The Jacobian of the constraints can be factored into two blocks:
+$Jc=\begin{bmatrix}Jc_{[0:m]}&Jc_{[m:N_{O}]}\end{bmatrix},\quad\text{with}\quad m=N_{O}-N_{C}$ (64)
+The mapping $\varphi_{x}$ represents the solution of a (non)linear system of equations, either explicitly or by using a root-finding solver (e.g., Newton method).
+According to Donti et al.
+[7], the differential of such mapping can be computed leveraging the implicit function theorem (as in OptNet [29]:
+$\frac{\partial\varphi_{x}}{\partial z}=J_{z}\varphi=-(Jc_{[m:N_{O}]})^{-1}Jc_{[0:m]}$ (65)
+Hence, the operator-norm bound leads to:
+$||J_{z}\varphi||\leq||(Jc_{[m:N_{O}]})^{-1}||\,||Jc_{[0:m]}||$ (66)
+$\left\|\varphi_{x1}-\varphi_{x2}\right\|\leq||J_{z}\varphi||\left\|z_{1}-z_{2}\right\|,$ (67)
+We observe that the Jacobian of the mapping (completion step) is not bounded, as it depends directly on the constraints.
+Indeed, while the second term $||Jc_{[0:m]}||$ can be tuned by re-scaling the variables in the neural network, there is no simple trick to condition the term $||(Jc_{[m:N_{O}]})^{-1}||$.
+The spectral norm of the inverse matrix can be approximated by the smallest singular value $\sigma_{min}(Jc_{[m:N_{O}]})$.
+Then, if the constraints Jacobian matrix block $Jc_{[m:N_{O}]}$ is not well-conditioned:
+$||(Jc_{[m:N_{O}]})^{-1}||\sim\frac{1}{\sigma_{min}}$ (68)
+$\sigma_{min}\rightarrow 0\,,\quad||J_{z}\varphi||\rightarrow\infty$
+Thus, we conclude that the completion step is in general not $1$-Lipschitz, with Lipschitz constant $\kappa=||(Jc_{[m:N_{O}]})^{-1}||$.
+This can potentially lead to instabilities during training and result in sub-accurate NNs, with respect to unconstrained counterparts (e.g., as in [28]):
+  • Adversarial robustness (or forward sensitivity): Small input perturbation can be amplified by a factor $\kappa$.
+  • Stability of gradient flow dynamics (or backward sensitivity): During backpropagation, according to [7], the upstream gradient can be computed as:
+    $\frac{d\ell}{dz}=\frac{\partial\ell}{\partial z}-\frac{\partial\ell}{\partial\varphi_{x}}\cdot\frac{\partial\varphi_{x}}{\partial z}$ (69)
+    We can bound the upstream loss as:
+    $\left\|\left\|\frac{\partial\ell}{\partial z}\right\|-\left\|\frac{\partial\ell}{\partial\varphi_{x}}\right\|\left\|\frac{\partial\varphi_{x}}{\partial z}\right\|\right\|\leq\left\|\frac{d\ell}{dz}\right\|\leq\left\|\frac{\partial\ell}{\partial z}\right\|+\left\|\frac{\partial\ell}{\partial\varphi_{x}}\right\|\left\|\frac{\partial\varphi_{x}}{\partial z}\right\|$ (70)
+    And considering the derived Lipschitz constant $\kappa$:
+    $\left\|\left\|\frac{\partial\ell}{\partial z}\right\|-\kappa\left\|\frac{\partial\ell}{\partial\varphi_{x}}\right\|\right\|\leq\left\|\frac{d\ell}{dz}\right\|\leq\left\|\frac{\partial\ell}{\partial z}\right\|+\kappa\left\|\frac{\partial\ell}{\partial\varphi_{x}}\right\|$ (71)
+    Hence, the gradients expand or shrink by a factor $\kappa$, potentially leading to undesired phenomena such as vanishing or exploding gradients.
+Appendix C Additional experiments
+C.1 Heuristic analysis of constrained learning and hyperparameters
+We report here additional heuristic observations and implications of our constrained learning routine on the simple function fitting case study.
+Specifically, (i) we detail the main results in terms of constraint feasibility, prediction accuracy, and inference time, (ii) we observe the positive influence of our approach on the training dynamics and loss convergence, and (ii) we study the impact of constrained learning hyperparameters such as $\epsilon_{T}$ and $\lambda_{D}$.
+Accuracy and constraint feasibility
 One of the major concerns in constrained learning is related to keeping the expressivity and approximation capabilities of NNs [59, 60] while enforcing constraints to be satisfied.
-Here, we evaluate the regression capabilities and robustness to spectral biases [53] of ENFORCE on oscillating functions ($f=5$) while assessing the average and maximum constraint deviation (i.e., residual).
-We compare the method with a traditional (unconstrained) multilayer perceptron (MLP) and a soft-constrained neural network sharing the same architecture.
-In the soft-constrained neural network, the constraint residual is added as an additional loss term balanced with a scalar weight ($\lambda_{y}$).
-We report here the outcomes of the best-performing soft-constrained network, with $\lambda_{y}=1$.
-We train all the NNs for 50,000 epochs, using Adam [54] optimizer with a learning rate of $0.001$.
 The main results are summarized in Table 1.
-ENFORCE outperforms the soft-constrained NN and the MLP by effectively minimizing the nonlinear constraint residual, guaranteeing arbitrary satisfaction with minor computational costs (Fig.
-3(c)).
-To quantify the relative impact, we normalized the residuals by scaling them with respect to the absolute maximum value of the variables within the considered range ($\max(|x|,|y_{1}|,|y_{2}|)$).
-It is worth noting how overall good performance (e.g., in NN) does not guarantee compliance with underlying model constraints, with deviations above 17%.
-By contrast, ENFORCE demonstrates a worst-case residual below 0.0001% (Table 1).
-The inference time for a batch of 1000 samples is 6 ms longer when using ENFORCE compared to an MLP.
-This amount should be regarded as additive, not multiplicative (e.g., 4x relative to the MLP).
+ENFORCE outperforms the soft-constrained neural network and the MLP by effectively minimizing the nonlinear constraint residual, guaranteeing arbitrary satisfaction with minor computational costs (Fig.
+2(c), Appendix C).
+To quantify the relative impact, we normalized the residuals by scaling them with respect to the absolute maximum value of the variables within the considered range.
+A good prediction accuracy (low MAPE) does not guarantee compliance with underlying constraints, with maximum relative deviations above 17%, i.e., in the MLP.
+By contrast, ENFORCE demonstrates a worst-case residual below the specified tolerance (Table 1).
+The inference time for a batch of 1,000 samples is 6 ms longer when using ENFORCE compared to an MLP.
+This amount should be regarded as additive (+6 ms), not multiplicative (e.g., 4x relative to the MLP).
 Indeed, the computational complexity is entirely attributed to the AdaNP module, meaning that the backbone architecture has no impact.
-Therefore, when applied to larger models (e.g., transformers), the relative computational impact becomes negligible.
-Indeed, the complexity is primarily influenced by the number of constraints and the batch size, with the latter being particularly significant when inference is performed on CPU hardware (see Section 4.1 and Appendix A.2).
+Therefore, if applied to larger backbones (e.g., transformers), the relative computational impact may become negligible.
 Besides constraint guarantee, ENFORCE also enhances the accuracy of the MLP (the $R^{2}$ increases from 0.994 to 0.999).
@@
 However, advancing deep learning methods to achieve near-perfect performance remains an open challenge in essential fields such as healthcare and self-driving cars, where even slight inaccuracies at the decimal level can lead to catastrophic consequences [10, 11].
-Interestingly, ENFORCE outperforms the MLP in data-scarce regions of the domain, which in this dataset correspond to the domain extremities (as shown in Fig.
-3(b)).
-More generally, ENFORCE also performs better under data-scarcity conditions when the models are trained on uniformly sampled fractions of the dataset (Fig.
-3(d)).
-This observation suggests that constrained learning may enhance data efficiency.
-Table 1: Regression accuracy and constraint guarantee of ENFORCE when compared with a multilayer perceptron (MLP) and a soft-constrained NN (Soft).
-Results for $\lambda=0.5$, $\epsilon_{T}=10^{-4}$, and $\epsilon_{I}=10^{-6}$ are reported.
-The inference on 1000 samples is run on a CPU (11th Gen Intel(R) Core(TM) i7-1185G7 @ 3.00GHz, 4 Core(s)) to prove efficiency also in the absence of GPU hardware.
-The inference time is also reported for a batch of 1000 predictions.
-MAPE [%] $R^{2}$ Avg Residual [%] Max Residual [%] Inference Time [s]
-MLP 0.339 $\pm$ 0.083 0.994 $\pm$ 0.003 1.47 $\pm$ 0.33 17.13 $\pm$ 3.94 0.002 $\pm$ 0.000
-Soft 0.944 $\pm$ 0.143 0.972 $\pm$ 0.002 1.55 $\pm$ 0.16 7.77 $\pm$ 0.40 0.002 $\pm$ 0.000
-ENFORCE 0.060 $\pm$ 0.028 0.999 $\pm$ 0.000 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.008 $\pm$ 0.003
+We extensively study the heuristic implications of our constrained learning routine on the illustrative function fitting, leveraging its simplicity (cf.
+Appendix C).
+Notably, we observe that ENFORCE outperforms the MLP even before projection (i.e., just using the backbone at inference).
 Effects of constrained learning
 Notably, ENFORCE outperforms the MLP even before the projection steps, demonstrating superior performance using only the neural network backbone (Fig.
-4(a), dashed-pink line).
+3(a), dashed-pink line).
 This can be attributed to the structure of the hard-constrained learning process, where the predictions are adjusted via projection to satisfy underlying constraints.
@@
 (d) Data Scarcity
-Figure 3: Prediction comparison between ENFORCE ($\lambda=0.5$, $\epsilon_{T}=10^{-4}$, $\epsilon_{I}=10^{-6}$) and a multilayer perceptron (MLP).
+Figure 2: Prediction comparison between ENFORCE ($\lambda_{D}=0.5$, $\epsilon_{T}=10^{-4}$, $\epsilon_{I}=10^{-6}$) and a multilayer perceptron (MLP).
 ENFORCE enhances the overall accuracy and guarantees satisfaction for highly nonlinear constraints.
 ENFORCE consistently performs better than a standard MLP even when trained on uniformly sampled fractions of the training dataset.
+Interestingly, ENFORCE outperforms the MLP in data-scarce regions of the domain, which in this dataset correspond to the domain extremities (as shown in Fig.
+2(b)).
+More generally, ENFORCE also performs better under data-scarcity conditions when the models are trained on uniformly sampled fractions of the dataset (Fig.
+2(d)).
+This observation suggests that constrained learning may enhance data efficiency.
 Training dynamics
 To understand the training dynamics of ENFORCE, we analyze the loss curves shown in Fig.
-4(a), where the training data loss of ENFORCE is compared to the MLP.
-Being interested in the effect of hard-constrained learning and to ease the visualization, we do not report here the loss curve of the soft-constrained NN.
+3(a), where the training data loss of ENFORCE is compared to the MLP.
+Being interested in the effect of hard-constrained learning and to ease the visualization, we do not report here the loss curve of the soft-constrained neural network.
 In this case study, AdaNP contributes to the learning process from the very early iterations (Fig.
-4(b), orange line), suggesting that the projection operations positively guide the optimization process.
-The combination of feasible predictions and minimization of projection displacement drives the learning process toward more optimal outcomes.
+3(b), orange line), suggesting that the projection operations positively guide the optimization process.
+The combination of approximated feasible predictions and minimization of projection displacement drives the learning process toward more optimal outcomes.
 The modified loss function effectively guides the training process toward smaller projection displacements (Fig.
-4(b), green dashed line).
+3(b), green dashed line).
 The displacement loss decreases consistently during training due to the influence of the penalty term in the loss function.
@@
 Furthermore, this decay in AdaNP depth is consistently observed across different training tolerance values, as illustrated in Fig.5(a).
-(a) Loss NN vs. ENFORCE
+(a) Loss neural network vs. ENFORCE
 (b) ENFORCE Losses and AdaNP
@@
 Constrained learning hyperparameters (training)
-We systematically analyze the influence of hyperparameters, such as the weighting factor $\lambda$ and the tolerance $\epsilon_{T}$, on the constrained learning process.
+We systematically analyze the influence of hyperparameters, such as the displacement weighting factor $\lambda_{D}$ and the tolerance $\epsilon_{T}$, on the constrained learning process.
 Fig.
@@
 Moreover, its influence is generally small compared to the variance of different training runs (Fig.
-5).
+4).
 Intuitively, a smaller tolerance $\epsilon_{T}$ necessitates deeper AdaNP modules, resulting in higher computational costs due to the increased number of neural projection layers.
 This effect is visible in Fig.
-6(a), where the depth of AdaNP during training is reported (i.e., number of projection layers).
+5(a), where the depth of AdaNP during training is reported (i.e., number of projection layers).
 The average depth of AdaNP increases to accommodate stricter tolerances.
@@
 More importantly, the required depth tends to have a slower decay during training, if compared to using less strict tolerances (as visible in Fig.
-6(a)).
+5(a)).
 Larger tolerances result, on average, in shallow AdaNP layers (approximately one layer).
@@
 Along with the minor impact on overall accuracy, this observation suggests setting the training tolerance $\epsilon_{T}$ to less stringent requirements.
-The regression accuracy is evidently affected by the choice of the loss weighting factor $\lambda$ (Fig.
-5).
+The regression accuracy is evidently affected by the choice of the displacement loss weighting factor $\lambda_{D}$ (Fig.
+4).
 Remarkably, unlike the challenging task of tuning weighting factors in soft-constrained methods [26], the constrained learning approach proposed here positively impacts accuracy regardless of the specific weighting factor chosen (as shown in Fig.
-5, the accuracy of ENFORCE is consistently greater than that of a standard MLP).
-However, an inappropriate choice of this parameter can result in suboptimal outcomes (e.g., when $\lambda=2$ in Fig.
-5).
+4, the accuracy of ENFORCE is consistently greater than that of a standard MLP).
+However, an inappropriate choice of this parameter can result in suboptimal outcomes (e.g., when $\lambda_{D}=2$ in Fig.
+4).
 Therefore, careful tuning of this hyperparameter is warranted.
@@
 Fig.
-6(b) illustrates the number of NP layers needed to satisfy the constraint under varying $\lambda$ and $\epsilon_{T}$.
+5(b) illustrates the number of NP layers needed to satisfy the constraint under varying $\lambda_{D}$ and $\epsilon_{T}$.
 The weighting factor is shown to reduce the required number of NP layers by half, with no additional cost during training.
@@
 This finding further supports the recommendation of employing shallow AdaNP modules during training, by relaxing the value of $\epsilon_{T}$.
-We conclude that the weighting factor $\lambda$ plays an important role by balancing the contribution of the projection displacement error.
-On the other hand, enforcing strict satisfaction during training at arbitrary small tolerance $\epsilon_{T}$ does not necessarily improve the overall outcome.
+We conclude that the displacement loss weighting factor $\lambda_{D}$ plays an important role by balancing the contribution of the projection displacement error.
+On the other hand, enforcing strict satisfaction during training with an arbitrary small tolerance $\epsilon_{T}$ does not necessarily improve the overall outcome.
 Figure 4: Influence of constrained learning hyperparameters on the accuracy of ENFORCE on the test set (note that here we plot the inverse of the mean squared error (MSE)).
-The weighting factor $\lambda$ favors the learning process if appropriately tuned.
+The weighting factor $\lambda_{D}$ favors the learning process if appropriately tuned.
 Conversely, the training tolerance $\epsilon_{T}$ exhibits a small impact on performance, suggesting it can be set based on available resources.
@@
 Figure 5: Dynamic evolution of AdaNP during training and inference when different training hyperparameters are chosen.
-At training time, AdaNP is deeper as a smaller constraints tolerance $\epsilon_{T}$ is chosen.
-6 Conclusions and outlook
-We developed ENFORCE, a method designed to guarantee neural network prediction satisfies any nonlinear constraint $c(x,y)=0$ within a specified tolerance $\epsilon$, without relying on external solvers or incurring prohibitive computational overhead.
-The effectiveness of the method is demonstrated in an illustrative case study chosen for its interpretability and ease of visualization.
-We found that (1) ENFORCE ensures the satisfaction of nonlinear constraints to within a tolerance $\epsilon$ for both average and maximum residuals, (2) the overall accuracy improves from traditional unconstrained MLP, achieving near-perfect performance on the test set, and (3) the constrained learning hyperparameters influence the learning process toward optimal performance; however, our method shows superior performance compared the unconstrained counterpart for a wide range of hyperparamer settings.
-Additionally, we observed promising performance in data-scarcity conditions (Fig.
-3(d)).
-This work opens several research avenues toward developing robust NNs that strictly adhere to underlying system knowledge.
-First, the current method can be extended to handle piecewise-defined constraints and nonlinear inequality constraints.
-Additionally, the requirement for the constraint to be a $\mathcal{C}^{1}$ function could be relaxed by leveraging sub-gradients [55].
-Finally, alternative (e.g., weighted) projection approaches could be explored to better account for the morphology and scaling of the constraints.
-The method has the potential to address specific challenges or complement existing approaches, including those based on neural networks and other machine learning models for regression tasks (e.g., Gaussian processes and support vector machines).
-For instance, hard constraints can be combined with soft-constraint techniques, such as PINNs, to reliably solve PDEs [56].
-Additionally, the method could enhance learning performance in partially annotated datasets by inferring missing information through available constraints.
-Finally, an interesting future direction could involve applying the AdaNP module to GenAI models, guiding the generation process toward domain-compliant samples, such as for synthetic data or image/video generation.
-Limitations
-In this paragraph.
-we highlight the main limitation of the proposed method.
-Firstly, the effectiveness of ENFORCE is highly dependent on the regression capabilities of the neural network backbone.
-When the model lacks sufficient complexity to achieve accurate predictions, ENFORCE provides limited benefit.
-Furthermore, the method becomes computationally intensive when applied to systems with a large number of constraints (e.g., more than 1000).
-This is due to the computational cost of each neural projection layer, which scales as $O(N^{3})$ because of the matrix inversion operation, where $N$ is the number of constraints.
-Further details are provided in Appendix A.2.
-Acknowledgements
-This research is supported by Shell Global Solutions International B.V., for which we express sincere gratitude.
-Appendix A Appendix
-A.1 Mathematical derivations
-This section provides the key mathematical derivations underlying the proposed method.
-These derivations are intended to illustrate the theoretical foundations of the approach and to support some of the discussions presented in the main text.
-Closed-form neural projection
-We derive here the closed-form expression defining a neural projection layer in Def. 1 (Section 4.1). Given the linearized projection problem in Eq. 3, we can define the Lagrangian function as:
-$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}\left(c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})\right)$ (6)
-Then, a local optimum can be found by solving the KKT conditions (i.e., primal and dual feasibility):
-$\nabla_{y}\mathcal{L}=(y-\hat{y})+\left.J^{T}_{y}c\right|_{x,\hat{y}}\lambda=0$ (7)
-$c(x,\hat{y})+\left.J_{y}c\right|_{x,\hat{y}}(y-\hat{y})=0$
-To simplify the notation, we define the linear system as:
-$(y-\hat{y})+B^{T}\lambda=0$ (8)
-$By-v=0$
-where:
-$B$ $=\left.J_{y}c\right|_{x,\hat{y}}\in\mathbb{R}^{N_{C}\times N_{O}}$ (9)
-$v$ $=\left.J_{y}c\right|_{x,\hat{y}}\hat{y}-c(x,\hat{y})\in\mathbb{R}^{N_{C}}$
-Solving the system, we obtain a closed form for the neural projection layer:
-$\tilde{y}=(I-B^{T}(BB^{T})^{-1}B)\hat{y}+B^{T}(BB^{T})^{-1}v$ (10)
-Deviation from Newton’s method
-To support the discussion raised in Section 4.1, we show how our method deviates from standard Newton’s method for solving nonlinear KKT conditions. Given a NLP:
-$\tilde{y}=$ $\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}$ (11)
-s.t. $\quad c(x,y)=0$
-With associated Lagrangian function:
-$\mathcal{L}(x,y,\lambda)=\frac{1}{2}(y-\hat{y})^{T}(y-\hat{y})+\lambda^{T}c(x,y)$ (12)
-The primal and dual feasibility can be derived as:
-$\nabla_{y}\mathcal{L}=(y-\hat{y})+J^{T}_{y}c(x,y)\lambda=0$ (13)
-$c(x,y)=0$
-If we linearized the system according to the Newton’s iteration, we obtain:
-$(y_{0}-\hat{y})+J^{T}_{y}c|_{y_{0},\lambda_{0}}\lambda_{0}+(y-y_{0})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{y_{0},\lambda_{0}}(y-y_{0})+\left.J^{T}_{y}c\right|_{y_{0},\lambda_{0}}(\lambda-\lambda_{0})=0$ (14)
-$c(x,y_{0})+J_{y}c|_{y_{0},\lambda_{0}}(y-y_{0})=0$
-Thus, assuming to center the linearization in the neural network prediction, i.e., $y_{0}=\hat{y}$, and choosing $\lambda_{0}=0$:
-$(y-\hat{y})+\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})+\left.J^{T}_{y}c\right|_{\hat{y},0}\lambda=0$ (15)
-$c(x,\hat{y})+\left.J_{y}c\right|_{\hat{y},0}(y-\hat{y})=0$
-We can conclude that, essentially, our NP layer solves a similar linear system (Eq.
-7) which does not comprise the term $\lambda^{T}\left.\mathbf{H}_{y}c\right|_{\hat{y},0}(y-\hat{y})$, hence avoiding the computation of the Hessian tensor $\mathbf{H}_{y}c$.
-A.2 Implementation details
-Batch local projection
+At training time, AdaNP is deeper as a smaller constraint tolerance $\epsilon_{T}$ is chosen.
+C.2 Scaling analysis for nonconvex problems with linear and nonlinear constraints
+In this section, we present detailed results from the scaling analysis conducted on the two classes of optimization problems presented in Section 5.2: (i) nonconvex problems with linear equality constraints, and (ii) nonconvex problems with nonlinear equality constraints.
+The following tables report key performance metrics of ENFORCE across varying problem sizes, including different numbers of constraints and optimization variables, and compare them with alternative deep learning-based methods and a traditional large-scale nonlinear programming solver such as IPOPT.
+Table 4 and Table 3 report the resulting metrics for the linearly constrained case and the nonlinearly constrained case, respectively.
+Table 4: Scaling experiments on a nonconvex optimization problem with linear equality constraints (Eq.
+9) evaluating performance across varying numbers of constraints ($N_{C}$) and variables ($N_{O}$).
+ENFORCE consistently predicts feasible and near-optimal solutions, outperforming alternative deep learning-based methods.
+DC3 is trained for a greater number of epochs than the other methods, until convergence is reached.
+Constraints ($N_{C}$) $50$ $70$ $150$
+Variables ($N_{O}$) $100$ $100$ $200$
+Method Metric
+IPOPT Obj. value $-11.11\pm 0.00$ $-4.84\pm 0.00$ $-10.64\pm 0.00$
+Max eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Mean eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Inf. time [s] $0.095\pm 0.033$ $0.13\pm 0.04$ $0.379\pm 0.060$
+Tr. time [min]
+Epochs
+MLP Obj. value $-27.43\pm 0.00$ $-27.43\pm 0.00$ $-52.99\pm 0.01$
+Max eq. $24.65\pm 0.08$ $24.89\pm 0.12$ $45.38\pm 0.56$
+Mean eq. $7.32\pm 0.00$ $7.19\pm 0.00$ $9.14\pm 0.02$
+Inf. time [s] $0.001\pm 0.000$ $0.001\pm 0.001$ $0.001\pm 0.001$
+Tr. time [min] $8.87\pm 0.18$ $8.89\pm 0.11$ $9.01\pm 0.25$
+Epochs $1000$ $1000$ $1000$
+Soft Obj. value $-10.10\pm 0.31$ $-1.86\pm 0.17$ $1.28\pm 0.32$
+($\lambda_{c}=5$) Max eq. $0.53\pm 0.04$ $0.79\pm 0.08$ $1.45\pm 0.43$
+Mean eq. $0.03\pm 0.00$ $0.06\pm 0.00$ $0.08\pm 0.00$
+Inf. time [s] $0.002\pm 0.000$ $0.001\pm 0.000$ $0.001\pm 0.000$
+Tr. time [min] $10.69\pm 0.51$ $10.72\pm 0.43$ $10.91\pm 0.46$
+Epochs $1000$ $1000$ $1000$
+Soft Obj. value $-10.69\pm 0.01$ $-4.18\pm 0.03$ $-8.18\pm 0.18$
+($\lambda_{c}=1$) Max eq. $0.54\pm 0.05$ $0.86\pm 0.05$ $1.47\pm 0.41$
+Mean eq. $0.05\pm 0.00$ $0.08\pm 0.00$ $0.09\pm 0.00$
+Inf. time [s] $0.001\pm 0.000$ $0.001\pm 0.000$ $0.001\pm 0.001$
+Tr. time [min] $10.69\pm 0.52$ $10.70\pm 0.46$ $10.85\pm 0.49$
+Epochs $1000$ $1000$ $1000$
+Soft Obj. value $-12.05\pm 0.00$ $-6.82\pm 0.01$ $-13.55\pm 0.02$
+($\lambda_{c}=0.1$) Max eq. $2.09\pm 0.03$ $2.51\pm 0.08$ $2.17\pm 0.12$
+Mean eq. $0.36\pm 0.00$ $0.43\pm 0.00$ $0.35\pm 0.00$
+Inf. time [s] $0.001\pm 0.001$ $0.001\pm 0.000$ $0.001\pm 0.001$
+Tr. time [min] $10.62\pm 0.61$ $10.73\pm 0.52$ $10.88\pm 0.54$
+Epochs $1000$ $1000$ $1000$
+DC3 Obj. value $-10.31\pm 10.07$ $-2.76\pm 0.06$ $-6.27\pm 0.07$
+Max eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Mean eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Inf. time [s] $0.003\pm 0.000$ $0.002\pm 0.000$ $0.004\pm 0.000$
+Tr. time [min] $22.96\pm 3.73$ $20.57\pm 8.30$ $25.18\pm 8.63$
+Epochs $3500$ $3500$ $3500$
+ENFORCE Obj. value $-11.50\pm 0.01$ $-4.86\pm 0.00$ $-10.59\pm 0.00$
+Max eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Mean eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Inf. time [s] $0.008\pm 0.001$ $0.010\pm 0.001$ $0.016\pm 0.002$
+Tr. time [min] $12.79\pm 0.03$ $12.72\pm 0.04$ $13.91\pm 0.07$
+Epochs $1000$ $1000$ $1000$
+Table 5: Scaling experiments on a nonconvex optimization problem with nonlinear equality constraints (Eq.
+10) evaluating performance across varying numbers of constraints ($N_{C}$) and variables ($N_{O}$).
+ENFORCE consistently predicts feasible and near-optimal solutions.
+In the simplest setting, it outperforms the nonlinear programming solver IPOPT in terms of solution quality.
+Constraints ($N_{C}$) $10$ $50$ $70$ $150$
+Variables ($N_{O}$) $100$ $100$ $100$ $200$
+Method Metric
+IPOPT Obj. value $-26.27\pm 0.00$ $-18.05\pm 0.00$ $-11.69\pm 0.00$ $-29.45\pm 0.00$
+Max Eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Avg. Eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Inf. time [s] $0.094\pm 0.032$ $0.268\pm 0.125$ $0.401\pm 0.167$ $3.40\pm 1.40$
+Tr. time [min]
+Epochs
+MLP Obj. value $-27.43\pm 0.00$ $-27.43\pm 0.00$ $-27.43\pm 0.00$ $-53.07\pm 0.00$
+Max eq. $214.95\pm 0.10$ $317.14\pm 0.01$ $317.97\pm 0.73$ $497.38\pm 4.64$
+Mean eq. $59.49\pm 0.03$ $69.63\pm 0.00$ $70.57\pm 0.01$ $118.39\pm 0.07$
+Inf. time [s] $0.001\pm 0.001$ $0.001\pm 0.001$ $0.001\pm 0.001$ $0.002\pm 0.001$
+Tr. time [min] $7.1\pm 0.3$ $10.1\pm 3.1$ $9.4\pm 0.3$ $10.6\pm 0.3$
+Epochs $1000$ $1000$ $1000$ $1000$
+Soft Obj. value $462.93\pm 29.91$ $>10^{5}$ $>10^{5}$ $>10^{5}$
+($\lambda_{c}=5$) Max eq. $16.31\pm 1.00$ $72.97\pm 0.74$ $73.63\pm 4.56$ $79.23\pm 3.81$
+Mean eq. $2.31\pm 0.10$ $16.81\pm 0.03$ $16.71\pm 0.03$ $16.72\pm 0.07$
+Inf. time [s] $0.001\pm 0.000$ $0.001\pm 0.001$ $0.002\pm 0.001$ $0.001\pm 0.001$
+Tr. time [min] $12.4\pm 0.1$ $12.8\pm 0.6$ $14.0\pm 0.6$ $15.1\pm 0.7$
+Epochs $1000$ $1000$ $1000$ $1000$
+Soft Obj. value $61.93\pm 3.31$ $>10^{4}$ $>10^{4}$ $>10^{4}$
+($\lambda_{c}=1$) Max eq. $15.67\pm 1.43$ $72.18\pm 2.21$ $73.80\pm 5.78$ $79.30\pm 3.70$
+Mean eq. $2.16\pm 0.06$ $16.81\pm 0.01$ $16.71\pm 0.04$ $16.68\pm 0.06$
+Inf. time [s] $0.001\pm 0.000$ $0.002\pm 0.001$ $0.001\pm 0.001$ $0.001\pm 0.000$
+Tr. time [min] $12.5\pm 0.1$ $11.0\pm 0.6$ $13.9\pm 0.6$ $15.0\pm 0.6$
+Epochs $1000$ $1000$ $1000$ $1000$
+Soft Obj. value $-18.29\pm 1.21$ $>10^{3}$ $>10^{3}$ $>10^{3}$
+($\lambda_{c}=0.1$) Max eq. $16.18\pm 0.55$ $76.56\pm 2.35$ $75.36\pm 2.70$ $78.35\pm 1.95$
+Mean eq. $2.05\pm 0.17$ $16.86\pm 0.05$ $16.77\pm 0.07$ $16.63\pm 0.01$
+Inf. time [s] $0.001\pm 0.000$ $0.002\pm 0.001$ $0.001\pm 0.001$ $0.002\pm 0.001$
+Tr. time [min] $11.5\pm 0.1$ $11.9\pm 0.1$ $13.9\pm 0.7$ $14.9\pm 0.8$
+Epochs $1000$ $1000$ $1000$ $1000$
+ENFORCE Obj. value $-26.37\pm 0.00$ $-16.68\pm 0.01$ $-7.75\pm 0.03$ $-27.77\pm 0.02$
+Max eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Mean eq. $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
+Inf. time [s] $0.013\pm 0.002$ $0.030\pm 0.005$ $0.049\pm 0.009$ $0.14\pm 0.08$
+Tr. time [min] $25.3\pm 0.1$ $35.8\pm 0.4$ $49.0\pm 0.9$ $69.4\pm 23.1$
+Epochs $1000$ $1000$ $1000$ $1000$
+Appendix D Implementation details
+D.1 AdaNP algorithm
+Algorithm 1 provides a high-level overview of the procedure underlying the AdaNP module.
+The depth of AdaNP (i.e., the number of projection iterations) adapts to satisfy the constraints requirements according to a specified tolerance.
+Algorithm 1 AdaNP: Adaptive-depth Neural Projection
+1:Input: input $x$, preliminary prediction $\hat{y}$, constraints $c$, tolerance $\varepsilon_{t}$, maximum depth $d_{\text{max}}$
+2:Initialize: depth counter $i\leftarrow 0$
+3:while $m(c(x,\hat{y}))>\varepsilon_{t}$ and $i<d_{\text{max}}$ do
+4:     Compute constraints Jacobian: $J_{y}c$
+5:     Compute: $B=\left.J_{y}c\right|_{x,\hat{y}}$
+6:     Compute: $v=B\hat{y}-c(x,\hat{y})$
+7:     Compute: $B^{*}=I-B^{T}(BB^{T})^{-1}B$
+8:     Compute: $v^{*}=B^{T}(BB^{T})^{-1}v$
+9:     Project: $\tilde{y}=B^{*}\hat{y}+v^{*}$
+10:     Update: $\hat{y}=\tilde{y}$
+11:     Increment: $i\leftarrow i+1$
+12:end while
+13:return $\tilde{y}$
+Here, $m(c(x,\hat{y}))$ represents some measure of the constraint residual, where $m(\cdot)$ can be the max or the mean operator.
+D.2 AdaNP activation algorithm
+In Section 4.4, we introduced a strategy for constrained learning during the early stages of training.
+The algorithm we propose (Algorithm 2) assesses the effectiveness of the projection operation by quantifying and comparing a task-specific loss measure ($m_{\ell}$) computed on both the preliminary and the projected predictions.
+This loss measure is different from the complete training loss function.
+In the presented experiments, denoting by $\bar{y}$ either $\hat{y}$ or $\tilde{y}$, depending on the context, we define:
+  • Function fitting: For regression tasks, $m_{\ell}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\bar{y}_{i}||^{2}$ is the standard mean squared error loss.
+  • Constrained optimization problem: In the context of unsupervised learning for parametric optimization problems, we define the loss measure as $m_{\ell}=\frac{1}{N}\sum_{i=1}^{N}f_{i}(x,\bar{y})+\frac{\lambda_{C}}{N}\sum_{i=1}^{N}||c(x_{i},\bar{y})||$.
+It is essential to include a penalty term for constraint violations, as strictly enforcing the constraints may inherently lead to larger values of the objective function.
+Algorithm 2 AdaNP activation algorithm during training
+1:Input: neural network $f_{\theta}$, input $x$, loss measure $m_{\ell}$
+2:Predict: $\hat{y}=f_{\theta}(x)$
+3:Project: $\tilde{y}=\mathcal{P}(\hat{y})$
+4:if $m_{\ell}(\tilde{y})>m_{\ell}(\hat{y})$ then
+5:     $\tilde{y}=\hat{y}$triangleright Discard the projection and use original prediction
+6:else
+7:     $\tilde{y}_{n}=\operatorname{AdaNP}(\tilde{y})$triangleright Activate AdaNP
+8:end if
+9:return $\tilde{y}_{n}$
+D.3 Batch local projection
 The computationally most expensive operation in the neural projection layer is the matrix inversion $(BB^{T})^{-1}$ (Def.
@@
 At inference time, $N=N_{C}$, since $B\in\mathbb{R}^{N_{C}\times N_{O}}$ and $BB^{T}\in\mathbb{R}^{N_{C}\times N_{C}}$.
-Hence, considering a number of constraint $N_{C}<10^{3}$, the matrix inversion is performed in $\sim 1$ million FLOPs, which is a negligible amount for most modern CPUs and GPUs.
+Hence, considering a number of constraints $N_{C}<10^{3}$, the matrix inversion is performed in less than 1 million FLOPs, which is an affordable amount for most modern CPUs and GPUs.
 During training, assuming the use of batch gradient descent and defining the batch size ($BS$) as the number of data points processed in a single iteration, an equivalent number of matrix inversions must be performed.
@@
 Modern deep learning libraries enable batch operations, such as matrix inversion, which reduce the effective complexity to $\mathcal{O}(N_{C}^{3})$ (i.e., scaling only with the number of constraints).
-Given these conditions and the capabilities of current hardware, the neural projection operation remains computationally efficient even during training when the number of constraints is fewer than a few thousand.
-This facilitates new computational approaches for handling and enforcing nonlinear constraints in neural networks.
+To invert the batch of matrices, we use the Cholesky factorization algorithm [39].
+Given these conditions and the capabilities of current hardware, the neural projection operation remains computationally efficient even during training when the number of constraints is in the order of a few hundred.
+Moreover, the complexity of this method is equivalent to other state-of-the-art methods such as DC3 [7], where each Newton’s step in the completion algorithm requires the inversion of a batch of $(N_{C}\times N_{C})$ matrices.
+D.4 Memory footprint
+For each neural projection layer, the AdaNP module creates four 3D tensors:
+  • Tensor $\mathbf{B}$ of shape $(BS,N_{C},N_{O})$
+  • Tensor $\mathbf{b}$ of shape $(BS,N_{C},1)$
+  • Tensor $\mathbf{B^{*}}$ of shape $(BS,N_{O},N_{O})$
+  • Tensor $\mathbf{b^{*}}$ of shape $(BS,N_{O},1)$
+Assuming 32-bit floating-point representation (i.e., 4 bytes per element), the memory required to store these tensors for a single projection step is:
+[…]
+Figures 6 and 7 illustrate how the memory requirement for a single projection layer is affected by variations in batch size, the number of predicted variables, and the number of constraints.
+The total memory usage of the unrolled AdaNP depends on the mode of operation:
+  • During training, tensors are retained at each of the $n$ projection steps (i.e., for gradient computation), resulting in:
+    $M_{\text{train}}=4\cdot n\cdot BS\cdot(N_{O}+1)\cdot(N_{C}+N_{O})$ (73)
+  • During inference, tensors are not retained across steps, and the peak memory usage is:
+    $M_{\text{infer}}=4\cdot BS\cdot(N_{O}+1)\cdot N_{O}$ (74)
+Figure 6: Memory footprint of a single projection layer $\mathcal{P}$.
+$N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output.
+The memory usage scales linearly with the batch size, with the growth rate determined by the number of predicted variables and constraints.
+For a batch size of 200, the memory requirement is on the order of tens of megabytes.
+Figure 7: Memory footprint for a fixed batch size of 200 samples.
+$N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output.
+For large-scale tasks involving thousands of variables and constraints, the AdaNP module can become memory-intensive, with each projection layer requiring more than 1 GB of memory.
+Figure 8: Memory usage during training and inference as a function of the number of projection layers $n$, for two problem scales.
+$N_{C}$ is the number of constraints, and $N_{O}$ is the dimension of the neural network output.
+Results are shown for a fixed batch size $BS=200$.
+Smaller-scale setup: $N_{C}=50$, $N_{O}=100$; larger-scale setup: $N_{C}=150$, $N_{O}=200$.
+Training memory increases linearly with $n$, with significantly higher usage in larger-scale problems.
+Inference memory remains nearly constant across all configurations.
+Appendix E Outlook and limitations
+This work opens several research avenues toward developing robust NNs that strictly adhere to underlying system knowledge.
+First, the current method can be extended to handle piecewise-defined constraints and nonlinear inequality constraints.
+Additionally, the requirement for the constraint to be a $\mathcal{C}^{1}$ function could be relaxed by leveraging sub-gradients [61].
+Finally, alternative (e.g., weighted) projection approaches could be explored to better account for the morphology and scaling of the constraints.
+The method has the potential to address specific challenges or complement existing approaches, including those based on NNs and other machine learning models (e.g., Gaussian processes and support vector machines).
+For instance, hard constraints can be combined with soft-constraint techniques, such as PINNs, to reliably solve PDEs [62].
+Additionally, the method could enhance learning performance in partially annotated datasets by inferring missing information through available constraints.
+Finally, an interesting future direction could involve applying the AdaNP module to GenAI models, guiding the generation process toward domain-compliant samples, such as for synthetic data or image/video generation.
+Limitations
+In this paragraph, we highlight the main limitation of the proposed method.
+Firstly, the effectiveness of ENFORCE is highly dependent on the regression capabilities of the neural network backbone.
+When the model lacks sufficient complexity to achieve accurate predictions, ENFORCE provides limited benefit.
+This is supported by the theoretical implications of the orthogonal projections on the constraints manifold discussed in this study.
+Furthermore, the method becomes computationally and memory-intensive when applied to systems with a large number of constraints (e.g., more than a few hundred).
+This is due to the computational cost of each neural projection layer, which scales as $O(N_{C}^{3})$ because of the matrix inversion operation, where $N_{C}$ is the number of constraints.
+Further details are provided in Appendix D.